### v1

# How to be an orthodox quantum mechanic

Abstract
This work sets out to answer a single question: what is the orthodox interpretation of quantum mechanics? However, we adopt a different approach to that normally used. Rather than carefully surveying the precise details of the thoughts of Bohr and Heisenberg, we extract an orthodoxy empirically. To do this we review a collection of $33$ textbooks on quantum mechanics, encompassing the most popular and prominent works of this nature. We then gauge their response to 12 propositions to build up a picture of exactly what is believed by an orthodox quantum mechanic. We demonstrate that this orthodoxy is largely unchanged over the past century, with some interesting emerging deviations, and has many aspects of Copenhagen-like viewpoints. However, it is more nuanced than some reductive characterisations that condense it down to the ontological primacy of the quantum state. The revealed orthodoxy has two main pillars: measurement inherently disturbs quantum states and these states refer to individual instances, not ensembles. More fully it entails that individual particles exist in wave-like super-positions and present particle behaviours only when forced to by outside influences. The act of measuring such a system inherently changes its state in a random fashion, manifesting in a form of measurement error that corresponds to the uncertainty principle. This implies that measurement does not reveal underlying values of quantum properties.
keywords:
Quantum foundations; Quantum interpretation
1 Introduction
How do scientists understand quantum mechanics? It is commonly asserted that the most popular interpretation of quantum mechanics (QM) is the “Copenhagen interpretation”. This is, however, notoriously vague [1, 2, 3, 4] and its characterisation tends to descend into hyperfine readings of Bohr and Heisenberg [2]. Rather than characterise this Copenhagen model, it seems more productive to instead determine what the consensus within the physics community actually is. In other words, we can look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
The question of what the orthodox view of quantum theory is has, of course, received previous attention in the literature. Home [4] argues that the primary characteristic of an orthodox viewpoint is that the state is framed as a complete description of a quantum system. Thus, this encapsulates both Copenhagen-like and ensembles-based theories. This latter inclusion stems from a claim that ensemble theories are internally incomplete unless they assume completeness of the quantum state (otherwise they must supply some hidden variables to assign definite values to measurable properties). Home refers to Feyeraband’s encapsulation of the standard position as being that quantum systems possess no definite properties over and above those that can be derived from the state itself [5]. Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [3]. These include the completeness of the state and the projection postulate. To produce something like a Copenhagen interpretation this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [6] described as “doing epistemology without ontology”. This all results in a similar view to Home, that it is the primacy of the quantum state that typifies the orthodox view. Interestingly, Stapp argues [2] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities. This seems somewhat difficult to square with the notion that the state is the most complete specification of the real situation, without some explanation of why additional information is inaccessible. These points reveal a tension in descriptions of a quantum orthodoxy: a Copenhagen-like view wants to claim a “completeness” or “finality” for QM (the “end of the road hypothesis” as Popper terms it [7]). However, it also contains a strong degree of instrumentalism, as acknowledged by both Home and Jammer. This latter aspect would deny any ontological aspects to the theory. It follows then that the state is merely a prescription for the prediction of experimental outcomes. It seems clear that a follower of an allegedly Copenhagen viewpoint could pick and choose between these two, somewhat opposed, strands of thought.
Clearly, it is not sufficient to characterise the orthodoxy of quantum interpretation by the single question “is the quantum state a complete description of reality?” At least not without presupposing the logical coherency of the orthodoxy and erasing the nuance in what we mean by “complete description”. This is typified by Home’s collapse of ensemble views into the orthodoxy. We must note that there is, so far, no experiment whose statistics are not accurately predicted by quantum mechanics. However, this does not itself entail all of the views associated with Bohr and Heisenberg that both Home and Jammer link to a standard viewpoint. What we need is a more detailed breakdown of interpretative beliefs that constitute the orthodoxy, divorced carefully from the common formalism.
How then do we study the beliefs of physicists as a community? The answer is that we turn to textbooks. The reason for this is that these works are presented for the education of a new generation of scientists. Thus, they are discussions of what we generally believe to be the ‘established facts’ of a given field. The key question is would the ‘established facts’ include an interpretive viewpoint? Perhaps textbooks function merely to convey a mathematical framework to their reader, with any interpretive remarks being negligible asides. This suggestion might have some validity, as a student of science can very often develop their intuition and understanding of a subject via that of the mathematics representing it. However, this is not straight-forwardly so in the realm of quantum mechanics. Evidenced by the proliferation of wildly different interpretations sharing the same (or very similar) mathematics. Thus, it is insufficient to pass on only a common mathematical core in this subject area. After all, most physics students will sketch out strikingly similar explanations of various unusual quantum results, even if their later work is done in deeply unrelated fields. They will also not merely bore a causal interlocutor with a barrage of equations. Where then would a student of quantum mechanics be exposed to an orthodoxy if not their textbooks? Perhaps in popular science media? However, one suspects that such media would itself be based on textbooks, rather than wade through the mire of competing interpretations in the active literature. An aspect of this can be seen in the fact that more recent textbooks devote more space to the interpretation of quantum phenomena (many have this as a separate chapter even). In addition, while working physicists may later deviate from the interpretations presented in reference works, we must acknowledge that their trajectory will be influenced by the initial conditions. Hence, the dominance of Copenhagen-like presentations in textbooks might lead to unorthodox interpretations which share many characteristics of this ‘reference orthodoxy’. An example of this would be the insistence that a quantum state does not apply merely to ensembles (as will be seen this is one of the most common assertions in textbooks). This is inherited by the vast majority of modern unorthodox interpretations (e.g. Bohmian mechanics [8], quantum Bayesianism [9], consistent histories [10], and relational quantum mechanics [11]), despite it not being necessary to the formalism [12]. Thus, even the beliefs of established researchers in quantum interpretation may be strongly correlated with the orthodoxy found in textbooks. It is therefore suggested that one cannot completely dismiss the interpretive remarks made in reference works on this subject. They fulfill a vital function in passing along a common understanding of quantum mechanics that is insufficiently conveyed by the mathematical formalism alone.
This work attempts an empirical investigation of the quantum orthodoxy via a survey of popular and widely used textbooks on quantum mechanics, for both graduates and undergraduates. This encompasses $33$ works, with dates of first publication spanning from 1930 to 2019. The sample starts with a set of works listed as “notable” on Wikipedia [^1] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics and expands it via textbooks that rank among Amazon best-sellers or are written by notable authors. To determine an orthodoxy empirically, we require some kind of framework of positions that might be part of the orthodoxy. Thus, we establish a set of 12 propositions, divided into three categories. The first of these encompasses how we understand the state in quantum mechanics. Here we identify three propositions that provide a Copenhagen-like framework: “the state is ontologically complete”, “the state does not merely describe an ensemble of identically prepared systems”, and “the state collapses upon measurement”. The second category revolves around the nature of particles with the propositions: “individual particles have wave-like properties”, “particle-like entities are described by a wave-particle duality”, “particles cannot simultaneously possess definite values of non-commuting observables”, “particles do not follow classical-like trajectories”, and “there is no adequate single picture of particle or wave”. Finally, we turn to the question of why quantum mechanics predicts statistics only. Here we have the propositions: “measurement inherently disturbs quantum systems”, “measurement creates the observed definite values”, “the uncertainty principle results from measurement disturbance”, and “the limitation to prediction of probabilities results from the uncertainty principle”. These will all be elaborated on in later sections of this paper. Each work is scrutinised and a trinary value is assigned to their agreement with each proposition (an in-between value is used when agreement can be strongly inferred from the context rather than explicitly stated). It can then be determined to what degree the works assent to each proposition on average and, by analysing the statistics, a set of propositions that robustly represent a consensus view point can be extracted (should one exist).
This paper is structured as follows: we first lay out the candidate propositions in section 2, then we detail the sample of textbooks considered in section 3, display example analyses of texts in section 4, and discuss the statistical properties in section 5. Finally, we draw conclusions in section 6.
2 Propositions of an orthodoxy
Defining all aspects of an interpretative orthodoxy is a difficult proposition. The approach taken here is to define three sub-categories of hopefully fundamental aspects of the interpretation of the quantum formalism, these being the nature of the quantum state, that of particles, and why our predictions are all statistical. Within each category we include propositions that source works can either agree with or dissent from. The selection of these questions was designed partly apriori and partly by assertions encountered in some of the works. We can then see why there is no mention of most of the vast multitude of interpretative approaches: they tend to feature only within scientific papers, rather than in textbooks designed to educate new generations of scientists. We reiterate that the choice of textbooks over research articles is made because we feel that the most distilled consensus is to be found in textbooks, which typically are formulated around what is considered “established knowledge” within a given field. It is entirely possible that this question set is incomplete (or contains some redundancy). However, we will see that the results deduced are quite statistically robust, even with the small sample size.
2.1 What is the quantum state?
The quantum state $\psi$ stands at the heart of the quantum formalism. Thus, it is an important interpretative task to decide what this mathematical object represents in the real world. The following propositions are designed to examine some common assertions to determine whether they form part of a quantum orthodoxy.
2.1.1 The state is ontologically complete
This proposition indicates that there is no further information about the ‘real state of affairs’ that can be obtained once one knows the quantum state of a given system. This stands opposed to cases like the de Broglie-Bohm interpretation, where the state is supplemented by hidden variables and a pilot wave equation. An ensemble interpretation might deny this proposition, but affirm that QM is statistically complete, not because it supplements the formalism or metaphysics, rather it leaves open that QM may be a quantum analogue of statistical mechanics. Another viewpoint that would disagree with this proposition is one that assigns only epistemic value to the state. For the purposes of this work, we will take a response such as “the state represents all we can know about a system” as affirming the proposition, unless they supply some explanation of why the full ontology is not accessible.
2.1.2 The state does not just describe ensembles
Here we mean that $\psi$ describes single realisations of a given system (shorthand “single instance $\psi$”). It is not limited to the statistics of an ensemble of similarly prepared systems. This means that $\psi$ represents the current state of any single run through a given experiment. Affirming the proposition requires that individual systems can exist in indefinite/superposition states.
2.1.3 The state collapses on measurement
The state is capable of superposition due to the linearity of the quantum formalism. This means that the state of a system can be a linear combination of mutually exclusive states. If this state is not merely epistemic, or defined over an ensemble, we need to explain why measurements only ever return a single eigenvalue. A very common means of doing so is that the state “collapses” from an indefinite superposition to a single definite eigenstate upon measurement.
2.2 The nature of particles
It is commonly asserted that particles exhibit wave behaviour in quantum mechanics. These propositions examine different formulations of these ideas to find out if any belong to an existing orthodoxy, as well as some related questions about particles in QM.
2.2.1 Particles are wave-like
Individual particles exhibit wave behaviours such as diffraction and interference. The waves do not merely appear in the statistics of quantum systems. Viewpoints that might deny this proposition would be a de Broglie-Bohm interpretation [8] (where wave and particle co-exist as linked entities) or an ensemble one (where the waves are merely statistical patterns).
2.2.2 Wave-particle duality is necessary
This is an assertion of the fundamental importance of wave-like behaviour, even if the wave-like behaviour is not displayed by individual particles. Unobserved particle-like entities behave like waves, or have wave-like probability densities, but like particles when we detect them. Thus, we have dual, mutually exclusive, descriptions of the same experiments. This can be viewed as a weaker form of the preceding proposition.
2.2.3 Complementarity
This proposition is usually specified as encompassing both the wave-particle duality and uncertainty principle. In this sense, it is a declaration that quantum systems have complementary properties that cannot be measured simultaneously. Importantly, this means there is no “single picture” that can be used to describe quantum systems, i.e. both wave and particle pictures are incomplete. It is this last statement that will be used to determine if a work supports complementarity. Even though it might be considered as being unnecessary, if it is given by two existing propositions, it is interesting to see if it is mentioned explicitly (as it sometimes presented as a key aspect of QM).
2.2.4 Particles do not have trajectories
There are no unambiguous trajectories followed by particles. These are an apparent phenomenon on the classical level only.
2.2.5 The uncertainty principle applies to single particles
A common joke about QM goes: Heisenberg is stopped by traffic police who ask him “Do you know how fast you were going?”. His reply is “No, but I know exactly where I am”. This is the essence of this proposition. That uncertainty applies to individual realisations of quantum systems. That is, if I measure the position of a particle I cannot also know its momentum simultaneously. More generally: systems cannot posses definite values of non-commuting observables simultaneously. This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble [7].
2.3 Why are our predictions statistical?
When we conduct quantum mechanical experiments we require measuring over an ensemble in order to compare to the probabilistic predictions of the theory. If our interpretation of the quantum state is not ensemble-based, we have to explain why this form of measurement should be the status quo. These propositions examine some explanations (and related ideas) that arise in the studied works, in order to determine which ones may belong to the orthodoxy.
2.3.1 Measurement inherently disturbs the system
Measurement inherently changes the state of a system while measuring it (except in the case of eigenstates). A viewpoint that denies this would note that non-interactive measurements exist in QM or that measurement has no general abstract definition.
2.3.2 Uncertainty results from disturbance
The source of the uncertainty principle is the fact that our measurements must disturb quantum systems. This results in a random statistical scatter in our measurements over an ensemble of similarly-prepared systems.
2.3.3 Uncertainty leads to probabilities
This proposition regards the uncertainty principle as a fundamental feature of nature. Particularly, this also requires single-particle uncertainty. The consequences of this is that measurements have probabilistic results, even though quantum mechanics is not formulated as fundamentally statistical. One might deny this if one recognises that the uncertainty principle is a consequence of quantum rules, rather than a postulate, or if one correctly notes the uncertainty principle is formulated in terms of statistical observables only.
2.3.4 Measurement creates observed values
The observed values of quantum properties are the result of an interaction between apparatus and system, in the sense that they do not reflect underlying pre-measurement values in cases that are not eigenstates of observable being measured. This stands in opposition to the usual scientific idea of measurement: that we are generally discovering the values of pre-existing properties. This assertion can take two forms: a strong form, where it is asserted that quantum systems do not actually possess definite properties till measured, and a weaker form. In the latter, authors might take the position that measurement merely changes pre-existing values. Both of these result in the pre-measurement value being indeterminate and irrelevant.
3 Sample of studied works
To perform this study we examine the $33$ popular textbooks listed in Table 1. These range in date of first publication from 1930 to 2019. This large timespan is of utility, as it allows us to determine whether there is any evolution within the consensus view of quantum mechanics. Additionally, we consider subdivision of these works into graduate ($N=14$) and undergraduate ($N=19$), as determined from the statements of the target audience in the work itself. These works were chosen to encompass the Wikipedia list on ‘Notable quantum mechanics textbooks’ [^2] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics ($N=17$) as well as to include as large a sample as feasible. To detect variation over time we also consider a sub-sample of works first published after 1980 ($N=24$). Note that Omnés work is included, even though it is a textbook that advocates against the orthodoxy in favour of “consistent histories”, as it provides its own description of the orthodox positions. In the case of [13], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the ‘conventional” approach. We will also make reference to a set of 4 works whom we do not include in any of the samples: [4, 14, 15, 12]. The first is excluded as it is a discussion of interpretations but will provide a useful point of comparison. The second two are analysed out of curiosity but are not targeted squarely as part of a physics syllabus. The final one will be analysed and found to be entirely outside the orthodoxy on almost every point of contention (so it is treated as an outlier).
Table 1: Sample of studied works. Note that the column ‘Year’ is that of first publication. ‘Level’ indicates undergraduate (U) or graduate (G). ‘Wiki’ is Y if the book is on the wikipedia list for notable textbooks in quantum mechanics. The four books at the bottom are surveyed but not included in various samples.
Title Author Year Ref Level Wiki
Basic quantum mechanics K. Tamvakis 2019 [16] U N
Lectures on quantum mechanics S. Weinberg 2013 [17] G N
The physics of quantum mechanics J. Binney & D. Skinner 2013 [18] U Y
Quantum concepts in physics M. Longair 2013 [19] G N
Quantum mechanics D. McIntyre et al 2012 [20] U Y
Quantum mechanics: a new introduction K. Konishi & G. Paffuti 2009 [21] U N
Introduction to quantum mechanics H. Müller-Kirsten 2006 [22] U Y
Quantum physics M. Le Bellac 2006 [23] G N
Introduction to quantum mechanics A. Philips 2003 [24] U N
Quantum Mechanics: Concepts and Applications N. Zettili 2001 [25] U Y
Understanding quantum mechanics R. Omnés 1999 [26] U N
Particles behave like waves T. Moore 1998 [27] U N
Introduction to quantum mechanics D. Griffiths 1995 [28] U Y
Quantum theory A. Peres 1993 [29] G Y
Quantum mechanics F. Mandl 1992 [30] U N
A modern approach to quantum mechanics J. Townsend 1992 [31] U Y
Quantum mechanics: an introduction W. Greiner 1989 [32] U N
Quantum mechanics A. Rae 1986 [13] U N
Modern quantum mechanics J. Sakurai 1985 [33] G Y
Principles of Quantum mechanics R. Shankar 1980 [34] G Y
An introduction to quantum physics A. French & E. Taylor 1978 [35] U Y
Quantum mechanics C. Cohen-Tannoudji 1977 [36] G Y
Quantum physics of atoms, molecules, R. Eisberg & R. Resnick 1974 [37] U Y
solids, nuclei and particles
Quantum physics S. Gasiorowicz 1974 [38] U Y
Feynman lectures on physics volume 3 R. Feynman 1965 [39] U Y
Elementary quantum mechanics P. Fong 1962 [40] U N
Introduction to quantum mechanics R. Dicke & J. Wittke 1960 [41] G N
Quantum mechanics A. Messiah 1959 [42] G N
Quantum mechanics: non-relativistic theory L. Landau & L. Lifshitz 1958 [43] G Y
Mathematical foundations of quantum mechanics J. von Neumann 1955 [44] G Y
Quantum theory D. Bohm 1951 [45] G N
Quantum mechanics L. Schiff 1949 [46] G N
The principles of quantum mechanics P. Dirac 1930 [47] G Y
Quantum mechanics: a modern development L. Ballentine 1998 [12] G N
Quantum computation and quantum information M. Nielsen & I. Chuang 2000 [15] G N
An introduction to quantum theory K. Hannabuss 1997 [14] G N
Conceptual foundations of quantum physics D. Home 1997 [4] - N
4 Analysis from example texts
To illustrate the analysis process we will consider several texts as examples. These are chosen because they all share a common case of having to infer agreement with at least one proposition (usually “measurement creates observed values”). This will allow the reader to determine that we are carefully evidencing degrees of assent or dissent in each case. Note that the sections below leave out propositions that are unaddressed by the given text.
4.1 Binney & Skinner
We begin with The physics of quantum mechanics by Binney & Skinner [18]. For “measurement disturbs systems” we find initial evidence on page 3: “Moreover, the act of measurement inevitably disturbs the atom, and leaves it in a different state from the one it was in before we made the measurement” (note that this is not in the context of the physics of a particular type of measurement). This is reinforced on page 130 in a discussion of the collapse hypothesis: “The Copenhagen interpretation does, however, contain a crucial insight into measurement by stressing that any measurement physically disturbs the system, so the system’s state after a measurement has been made is different from what it was earlier”. For “uncertainty linked to disturbance” we find evidence for endorsement on page 3 (“We rarely know the exact state that our measuring instrument is in before we bring it into contact with the system we have measured, so the result of the measurement of the atom would be uncertain even if we knew the precise state that the atom was in before we measured it, which of course we do not”), page 28 (a discussion of the uncertainty principle in terms of measurement uncertainty) which is subsequently reinforced on pages 130 and 131. For single particle uncertainty we find on page 28 “the position of a particle that has well defined momentum is maximally uncertain” as evidence for adherence to this position. For “particles are wave-like” we find a discussion opposing this on page 7, particularly “we cannot consistently infer from this correspondence that particles are manifestations of waves”. For “collapse”, the work opposes this idea explicitly stating collapse is not physical on page 15 and that the hypothesis is false on page 130. For “particle-wave duality” the authors never mention the concept explicitly but already oppose “single particles are wave-like” outside of quantum field theory contexts (page 7). For “non-ensemble $\psi$” we find on page 14 “in quantum mechanics an individual photon has an amplitude to be linearly polarised in a any chosen direction and an amplitude to be circularly polarised in a given sense”, so states correspond to individual system realisations. This is reinforced on page 71 where they state “In quantum mechanics, complete information about any system is contained in its ket $\left|\psi\right>$”. For “probability from uncertainty” the authors dissent, as we have seen that they attribute probability to the uncertain state of the measuring apparatus and its subsequent disturbing of the system. For “measurement creates observed values” we infer assent, their view seems to be that systems are disturbed by measurement, changing the values found rather than creating them altogether. We have already seen that the authors view the state as ontologically complete, an “extended” theory is possible, but it merely accounts for the actual physics of measurement (page 130 and 131).
4.2 Tamvakis
Now we consider Basic quantum mechanics by Tamvakis [16]. On page 86 it is outlined, as part of the axioms of quantum mechanics, that the state changes during measurement indicating support for “measurement disturbs systems”. On page 5 uncertainty is described in terms of single particle measurement precision or “if we make a measurement on any object and determine its position along a given direction with uncertainty $\Delta x$, its momentum along this direction will necessarily be known with uncertainty no less than $\frac{h}{4\pi\Delta x}$”, an indication of “single particle uncertainty”. On page 6 the de Broglie relation is discussed as the wave aspect of a single particle, implying individual particles are wave-like. Pages 86 and 87 discuss a non-unitary evolution during measurement and call this collapse, this re-iterated on page 442. On page 5: “the particle aspects of a system are complementary to its wave aspects” directly supports complementarity. On page 7 it is stated that “the wave function of the particle characterises the system fully at each point in time”. This indicates ontological completeness and a single realisation/particle association also supported by the previous quote. Page 5 describes the uncertainty principle in terms of knowledge of $x$ and $p_{x}$ “thus, viewed from a classical viewpoint, in a microscopic system, complete knowledge of $x$ and $p_{x}$ is impossible”, we infer this is indicating that probabilities result from incomplete knowledge implied by the uncertainty principle. For “measurement observed creates values” we note that it is stated on page 86 that a superposition represents the pre-measurement state, since this cannot have a definite property value we must conclude that definite property values are created, or at least changed, during measurement. This will satisfy an assent to “measurement observed creates values” as a pre-measurement state with definite values cannot be recovered as the author describes measurement as “acausal” on page 87.
4.3 Zettili
Next we consider Quantum Mechanics: Concepts and Applications by Zettili [25]. On page 24 “measurement interferes with the state of microscopic objects” and “it is the effects of interference by the equipment on the system which is the essence of quantum mechanics” on page 165 demonstrate support for measurement disturbance. On page 169 uncertainty is described as a result of measurement “perturbing” the system. “Single particle uncertainty” is affirmed on page 28 by “If the $x$ component of the momentum of a particle is measured with an uncertainty $\Delta p_{x}$, then its $x$ position cannot, at the same time, be measured more accurately than $\Delta x=\frac{\hbar}{2\Delta p_{x}}$”. Support for single particles being wave-like is given on page 18: “each material particle behaves as a group of waves”. Pages 24 and 25 indicate “we cannot follow a microscopic particle along its motion, nor can we determine its path”, while 51 discusses wave packet spreading only meaning the “position cannot be known exactly”. These are taken to indicate trajectories exist but are unknown. On page 158 collapse is included in the postulates of quantum mechanics. On page 26 “Microscopic systems, therefore, are neither pure particle nor pure waves, they are both” as well as direct invocations on the same page, suffices for duality and complementarity. For ontological completeness, page 158 asserts “the state contains all the needed information about the system”. On page 30 we find “In quantum mechanics the state (or one of the states) of a particle is described by a wave function $\psi(r,t)$ corresponding to the de Broglie wave of this particle; so $\psi(r,t)$ describes the wave properties of a particle” indicates that the state applies to single instances. We combine previous statements from 165 and 169 to conclude that probability is introduced by measurement disturbance (uncertainty). Based on the axioms from 158 we infer that measurement creates observed values, as pre-measurement states can be superpositions and apply to individual particles/realisations.
4.4 Mcintyre, Manogue, & Tate
We now turn to Quantum mechanics by McIntyre et al [20]. On page 8 we find that measurements disturbing systems is a “key feature” of quantum mechanics. On page 59, in the context of the uncertainty principle, it is stated that “the lack of ability to measure all spin components simultaneously implies that the spin does not really point in a given direction”. This indicates the uncertainty principle is being applied to individual particles. Combining this with similar discussion on 208 (about position and momentum) implies a link between uncertainty and disturbance. A confusing discussion on 207 about the double-slit seems to suggest both that a wave picture is just for finding particle probabilities but also that the waves have to actually interfere (passing through two slits). The same discussion refers to Feynman and Cohen-Tannoudji which seems to indicate the ultimate outcome is individual particles are wave-like too. Page 192 indicates we cannot know a particles trajectory but does not deny its existence. Page 48 includes collapse in the postulates. Wave-particle duality is explicitly invoked as a correct description of quantum systems on pages 121 and 207. Complementarity in terms of position and momentum is discussed on pages 192 and broadened to particle and wave properties in 207 and 208. “The state of quantum mechanical system, including all the information you can know about it, is represented mathematically by a normalised ket $\left|\psi\right>$” covers ontological completeness. Discussion of Schrödinger’s cat on 167 and 168 is illuminating in terms of the role of ensembles. Here it is asserted that only decoherence prevents superpositions at classical scales and interference is used to ascribe physical signficance to superposition, beyond mere statistics. Thus, we see that the authors take superpositions to describe the state of individual realisations. Acceptance of real superpositions combined with the stated quantum postulates requires definite values are created by measurement.
4.5 Konishi & Paffuti
Finally we review Quantum mechanics: a new introduction by Konishi & Paffuti [21]. On page 29 the authors support the notion that all measurement causes disturbance by stating that “the disturbance caused by the process of measurement on the microscopic system under examination cannot be controlled beyond a certain limit.” The unqualified illustration of the uncertainty principle via “Heisenberg microscopes” on page 28 indicates the authors link uncertainty to measurement disturbance. On page 26 we find “the classical concept of a particle trajectory no longer makes sense. An electron cannot simultaneously have a definite position and momentum.” This demonstrates both that the uncertainty principle is applied to individual particles and that the authors reject trajectories for quantum objects. This comes with a caveat on page 29 that uncertainty principle is not itself fundamental, but derives from the “dynamical properties of the physical processes involved in the measurement.” On page 6 the authors indicate that single particles are wave-like by asserting that wave-particle duality “concerns the properties of an individual electron” and not a population. This, combined with page 28 highlighting the importance of wave-particle duality to uncertainty relations, supports a more general duality thesis too. On pages 31 and 32 we find that the authors believe state reduction is “necessary for the theory to make sense at all.” This is furthered on 518 and 519 where they argue collapse is experimentally confirmed. On page 25 the authors describe a superposed polarisation state as “the state of a single photon”. Additionally, on page 23 the authors ascribe “physical properties” to superposition states that are “somewhere between” those of the constituents. These indicate that quantum states do not apply to ensembles but to single realisations of systems. On page 29 the authors agree that probability derives from the uncertainty principle “the probabilistic nature of quantum mechanical pediction for the results of measurement is an inevitable consequence of Heisenberg’s uncertainty principle.” Although the discussion on 518 and 519 is ambiguous as to whether values are being revealed or created (the authors use “indefinite” properties of a superposition here to mean “unpredictable”), we must infer that they are at least being altered by measurement to comport with previous assertions. On page 21 the authors state that “knowledge of the wavefunction amounts to complete knowledge of the state” which, combined with prior assertions fully supports ontological completeness.
5 Results and discussion
In Table 2 we display all the data used for this study. Each work’s response to the propositions is characterised by a degree of agreement. This is kept as simple as possible with three principle values: $1$ indicates agreement, $-1$ dissent, and 00 as not mentioned. We expand this via either of ±$\pm 0.5$ to mean ‘inferred for logical consistency’. Inference of agreement/dissent is only made if the case is strong (in order to allow for a conservative analysis). We proceed to study the responses by observing the frequency of agreement. Note that this computation counts ±$\pm 0.5$ responses as half a count when determining frequencies. Our raw data keeps the page references separate for each proposition (available on request).
[…]
Tamvakis [16] 1 0 1 1 0 1 1 1 1 1 0.5 1 5,6,7,86,87,442
Weinberg [17] 0 0 0.5 -1 0 -1 0 0 1 0 0 -1 21,25,82,83,88,95,97
Binney & Skinner [18] 1 1 1 -1 0 -1 -1 0 1 -1 0.5 1 3,7,8,14,15,28,71,130,131
Longair [19] 1 1 1 1 0.5 1 1 1 1 1 0.5 1 178,183,359,360,363,367
Mcintyre et al [20] 1 0.5 1 1 -1 1 1 1 1 0 0.5 1 7,8,48,59,121,167,168,192,207,208
Konishi & Paffuti [21] 1 1 1 1 1 1 1 0 1 1 0.5 1 6,21,23,25,26,28,29,518,519
Müller-Kirsten [22] 1 0 1 1 0 0 1 1 1 1 1 1 9,13,17,19,20,169
Le Bellac [23] 1 0 1 1 0 1 1 0 1 0 1 1 22,24,97,101,102,104,179,185
Phillips [24] 1 1 1 1 1 1 1 0 1 0 1 -1 6,11,13,16,21,27,50
Zettili [25] 1 1 1 1 -0.5 1 1 1 1 0.5 0.5 1 18,24,25,26,28,30,51,158,165,169
Omnés [26] 1 1 1 1 1 1 1 1 1 1 1 1 31,45,47,49,53,91
Moore [27] 1 0 -1 1 0.5 1 1 1 1 0 1 -1 69,77,81,106,107,108,135
Griffiths [28] 1 1 -1 0 0.5 1 1 0 1 -1 1 1 1,2,3,4,5,19,112,420,431
Peres [29] -1 0 -1 1 -1 0 1 1 -1 -1 1 1 5,11,14,20,24,25,26,27,93
Townsend [31] 1 1 1 1 0 1 0 0 1 0 1 1 6,8,10,13,16,25,165,211
Mandl [30] 1 1 0.5 1 0 0 0 0 0.5 0 0 1 3,50,83
Greiner [32] 1 0 1 1 1 1 1 0.5 1 0 1 1 29,34,38,51,53,469,473
Rae [13] 0.5 -1 -1 1 1 1 1 1 1 0 1 1 6,8,10,61,73,79,261,273,287
Sakurai [33] 1 1 -1 -1 0 1 0 0 1 0 0 1 6,11,23,24,101
Shankar [34] 1 1 1 1 1 1 1 0 1 0 0.5 1 111,112,113,115,116,129,140
French [35] 1 1 1 1 0.5 1 1 0 1 0 0.5 0 72,76,81,85,92,93,243,317,328
Cohen-Tannoudji [36] 1 1 1 1 1 1 1 1 1 0 0.5 1 14,19,21,27,28,221,252,253,279
Eisberg & Resnick [37] 1 1 1 1 0 0 1 1 1 1 0.5 0 64,65,68,70,73,74,76,87,88,147
Gasiorowicz [38] 1 1 1 1 0 0.5 1 0 1 0 1 1 13,24,32,94,318,323,327
Feynman [39] 1 1 1 1 1 1 1 1 1 1 1 1 1-(6,7,9,10,11),2-(1,2,3,8,10)
Fong [40] 1 1 1 1 0 1 1 1 1 1 1 1 28,32,50,51,52,53,114,116,123,124
Dicke & Wittke [41] 1 1 1 1 0 1 1 1 1 0 1 1 14,15,16,24,27,40,74,91,110,115,120
Messiah [42] 1 1 1 1 1 1 1 1 1 0 0.5 1 20,21,49,116,119,140,143,144,153,157,162
Landau & Lifshitz [43] 1 1 1 1 1 1 0 0 1 1 1 1 2,3,4,5,6,24,47,157
von Neumann [44] 1 0 1 0 1 1 -1 0 1 1 1 1 4,210,214,216,247,326,403
Bohm [45] 1 1 1 1 1 1 1 1 1 1 0.5 1 59,73,80,99,104,117,120,145,160
Schiff [46] 1 1 1 1 1 0 1 1 1 0 0.5 1 3,6,7,8,13,16,21
Dirac [47] 1 1 1 1 0 1 1 0 1 0 1 1 4,7,8,9,16,34,46,52,98
In Figure 1, we see the agreement frequencies for each of the questions in turn. We display these in terms of 4 samples: the full sample (black), undergraduate books (yellow), graduate texts (blue), works on the ‘Notable books’ list from Wikipedia (red), and ‘modern’ works (gray). Full information on the degree of agreement, and the $p$ value computed from a binomial distribution are included in Tables 3 to 7. These $p$ values are drawn from the binomial distribution and represent the probability of the given degree of assent to arise by chance given the size of the sample. We choose a random binomial distribution because we consider a zero consensus result would be equal frequencies of assent ($1$) and non-assent (where we group 00 and $-1$ together). The grouping is done because no assertion in the sample has a consensus answer of 00 or $-1$. Thus, we take a high $1$ frequency as agreement with this assertion is “in the orthodoxy”. We also include a $\sigma$ value, the number of standard deviations of the binomial distribution that the assent frequency corresponds to.
Figure 1: Mode frequency of the responses to each question. The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), and modern (grey).
Table 3: Degree of consensus from the full sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.77 1.62e-04 3.1
Non-ensemble psi 0.95 1.16e-10 5.2
Ontologically complete 0.85 5.46e-06 4.0
No trajectories 0.45 3.64e-01 0.5
Complementarity 0.53 2.43e-01 0.3
Wave-particle duality 0.79 1.62e-04 3.3
Single particle uncertainty 0.82 3.31e-05 3.7
Single particle wave-like 0.85 5.46e-06 4.0
Measurement creates values 0.71 2.28e-03 2.4
Probability from uncertainty 0.35 8.14e-02 1.7
Measurement disturbs system 0.92 6.54e-08 4.9
Uncertainty linked to disturbance 0.71 2.28e-03 2.4
Table 4: Degree of consensus from the undergraduate sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.76 9.61e-03 2.3
Non-ensemble psi 0.97 1.91e-06 4.1
Ontologically complete 0.79 2.21e-03 2.5
No trajectories 0.39 3.24e-01 0.9
Complementarity 0.55 1.80e-01 0.5
Wave-particle duality 0.84 3.64e-04 3.0
Single particle uncertainty 0.82 3.64e-04 2.8
Single particle wave-like 0.89 3.81e-05 3.4
Measurement creates values 0.76 9.61e-03 2.3
Probability from uncertainty 0.39 3.24e-01 0.9
Measurement disturbs system 0.97 1.91e-06 4.1
Uncertainty linked to disturbance 0.71 9.61e-03 1.8
Table 5: Degree of consensus from the graduate sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.79 6.47e-03 2.1
Non-ensemble psi 0.93 6.10e-05 3.2
Ontologically complete 0.93 6.10e-05 3.2
No trajectories 0.54 2.12e-01 0.3
Complementarity 0.50 3.95e-01 0.0
Wave-particle duality 0.71 2.87e-02 1.6
Single particle uncertainty 0.82 9.16e-04 2.4
Single particle wave-like 0.79 6.47e-03 2.1
Measurement creates values 0.64 8.98e-02 1.1
Probability from uncertainty 0.29 8.98e-02 1.6
Measurement disturbs system 0.86 9.16e-04 2.7
Uncertainty linked to disturbance 0.71 2.87e-02 1.6
Table 6: Degree of consensus from the ‘Wikipedia notable’ sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.74 2.45e-02 1.9
Non-ensemble psi 0.94 7.63e-06 3.6
Ontologically complete 0.88 1.37e-04 3.2
No trajectories 0.35 1.66e-01 1.2
Complementarity 0.41 3.15e-01 0.7
Wave-particle duality 0.71 2.45e-02 1.7
Single particle uncertainty 0.82 1.17e-03 2.7
Single particle wave-like 0.76 6.36e-03 2.2
Measurement creates values 0.74 2.45e-02 1.9
Probability from uncertainty 0.32 1.66e-01 1.5
Measurement disturbs system 0.94 7.63e-06 3.6
Uncertainty linked to disturbance 0.79 1.17e-03 2.4
Table 7: Degree of consensus from the ‘modern’ sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
Assertion Affirmation $p$ $\sigma$
Collapse 0.71 1.13e-02 2.0
Non-ensemble psi 0.90 1.49e-06 3.9
Ontologically complete 0.79 7.72e-04 2.9
No trajectories 0.35 7.58e-02 1.4
Complementarity 0.40 2.71e-01 1.0
Wave-particle duality 0.71 1.13e-02 2.0
Single particle uncertainty 0.60 1.54e-01 1.0
Single particle wave-like 0.75 3.31e-03 2.4
Measurement creates values 0.67 3.20e-02 1.6
Probability from uncertainty 0.23 1.13e-02 2.7
Measurement disturbs system 0.85 1.39e-04 3.5
Uncertainty linked to disturbance 0.48 5.81e-01 0.2
We begin with propositions related to the nature of the quantum state. Interestingly, we see that not all the books regard $\psi$ as ontologically complete. However, the support for this proposition is robust with assent frequency $0.85$ ($p=5.46\times 10^{-6}$) in the full sample, with all of sub-samples having assent frequency $\geq 0.79$ ($p\leq 7.72\times 10^{-4}$). It is interesting that the lowest support for this occurs within the undergraduate and ‘modern’ samples. In terms of dissent, Phillips [24] and Moore [27] limit themselves to saying that $\psi$ can determine all the probabilities associated with a system but do not go on to say this is all the information allowed. Whereas, [17] openly suggests that interpretive difficulties might be overcome by finding a more general theory to supersede quantum mechanics. We also see strong support (a minimum frequency of $0.9$ with $p\leq 6.1\times 10^{-5}$) for the notion that a quantum state describes a single instance (as opposed to an ensemble). Notably, this achieves $0.95$ ($p=1.16\times 10^{-10}$) assent in the full sample and the undergraduate works are near unanimous on this count. Support is weaker for state collapse, where we obtain sub-sample assent frequencies above $0.71$ ($p\leq 2.45\times 10^{-2}$) while achieving $f=0.77$ ($p=1.62\times 10^{-4}$) in the full sample. From these results we can also see that an orthodox view requires the state to contain the entire ontological situation and apply to individual particles. This, in turn, requires us to believe that superposition states can be occupied by individual systems. Of the three propositions, only affirmation of collapse reduces slightly if we consider sources after 1980. Thus, the orthodoxy on these issues seems largely unchanged over the last century. It is noteworthy that, in a sample of $33$ works, there are only two dissenters to the notion that quantum states collapse: Weinberg [17] as well as Binney & Skinner [18]. Additionally, the topic is not mentioned by Griffiths, Peres, Müller-Kirsten, and Schiff [28, 29, 22, 46]. Similarly, only Peres [29] dissents from the notion that quantum states describe individual instances. Interestingly, we note that Home [4] argues that collapse is not essential to characterise the orthodoxy. However, this is strongly disputed by the data gathered here. Overall, we see that the orthodoxy has it that the quantum state is ontologically complete, refers to individual instances of systems, and collapses on measurement.
Now we turn to propositions on the nature of particles. A strong degree of consensus is clear around the notion that individual particles display wave-like behaviour with assent frequency $0.85$ at ($p\leq 5.46\times 10^{-6}$), sub-sample frequency $\geq 0.75$ ($p\leq 6.36\times 10^{-3}$). The only dissenters here being Weinberg [17], Binney & Skinner [18], and Sakurai [33]. The topic is also not discussed in von Neumann [44] (perhaps unsurprising given the mathematical focus) and Griffiths [28].
Reasonable support also exists for the notion of wave-particle duality with assent $0.79$ ($p\leq 1.62\times 10^{-4}$) in the full sample, sub-sample frequency $\geq 0.71$ ($p\leq 2.87\times 10^{-2}$). Here only von Neumann [44] and Binney & Skinner [18] outright denying this proposition. It is also not mentioned by Weinberg [17], Mandl [30], Townsend [31], and Sakurai [33].
For the proposition that individual particles being subject to the uncertainty principle, that is, one cannot know a given particle’s position and momentum (for example) simultaneously to arbitrary precision, we see more controversy. This is so because the full sample finds assent of $0.82$ ($p=3.31\times 10^{-5}$), however, the ‘modern’ subs-sample only registers $0.6$ ($p=0.154$). The reason for this change is that Griffiths [28], Peres [29], McIntyre [20], Moore [27], and Sakurai [33] all actively deny this proposition. This shows some evidence of a change in the consensus over time. However, a new orthodoxy is not yet entrenched, as Binney & Skinner [18], Longair [19], Townsend [31], Konishi & Paffuti [21], Omnés [26], Shankar [34], Phillips [24], Greiner [32], and Le Bellac [23] all affirm the proposition. Additionally, it is reasonably inferred that Mandl [30] and Weinberg [17] also support it. Home [4] does not include this in their conception of the orthodoxy on the grounds that is a flawed proposition. As argued in [7], the uncertainty principle could not function in the suggested manner. If the measurements from individual particles had the uncertainty principle’s scatter, the final ensemble distribution would have a larger uncertainty. Thus, preventing the verification of the principle as derived, where it applies to standard deviations over an ensemble of trials. Despite scepticism on this point, it is clearly wrong to exclude this from the orthodoxy, as the idea is uniformly affirmed by older works and still prevalent, but less so, in modern texts.
It is notable that affirmation of complementarity is generally weak, even the full sample only finds an agreement frequency of $0.53$ ($p=0.243$). This is not due to denial but simply it never being mentioned in many later texts. Regardless, it cannot have a place in any robust orthodoxy. We also see that the issue of whether particles possess trajectories is contentious, with frequency $0.45$ ($p=0.364$) in the full sample. This becomes even more contentious in the ‘modern’ sample with $0.35$ ($p=0.0758$). This proposition still finds agreement in modern works like Konishi & Paffuti [21], Cohen-Tannoudji [36], Shankar [34], Greiner [32], and Phillips [24], with partial agreement in Moore [27] and Longair [19]. However, the mode response is 00 in the undergraduate, ‘Wikipedia notable’ and ‘modern’ samples. So there is some indication that this proposition is increasingly marginalised by the orthodoxy. It might also follow that those who affirm the “single particle uncertainty” proposition should, for consistency, agree that particles have no trajectory (at least for resolutions finer than that allowed by uncertainty). However, we have tried to avoid forcing logical consistency on authors where there is some ambiguity or nuance.
Thus, we see that the orthodoxy holds that particles are individually wave-like, have a dual wave-particle nature reflected in their statistics, and that the uncertainty principle applies to measurements performed on single instances of systems (rather than acting only as a statistical deviation over an ensemble).
Next, we turn to propositions around the question of why quantum mechanics predicts only statistical results. There is clear consensus, regardless of whether or not we confine our attention to modern works, that measurement disturbs quantum systems with assent $0.92$ ($p=6.54\times 10^{-8}$) in the full sample. Notably, the undergraduate sample has a frequency of $\sim 1.0$ for this proposition and only the graduate and ‘modern’ samples drop below $0.94$. Indeed, only Peres [29] actively dissents. Their dissent is justified by the possibility of non-interactive measurement in QM [48]. There is also support for the idea that measurement itself actually creates the observed values with frequency $0.71$ ($p=2.28\times 10^{-3}$) in the full sample while retaining $0.67$ ($p\leq 3.2\times 10^{-2}$) in the ‘modern’ sub-sample. It should be noted that the affirmation is only stated openly 16 out of 29 times, despite being a plausible corollary to “measurement disturbs quantum systems” and that there are no active dissenters, not even [29]. The low significance thus represents the number of times the assent was only inferred. Only three authors do not provide grounds for inferring agreement: Sakurai [33], Mandl [30], and Weinberg [17]. The common case for inference is as follows: $\psi$ applies to individual systems, hence superpositions are real. Measurement disturbs the system and changes a superposition to an eigenstate, so it is either changing an unknown underlying value or creating a definite one where there was none. In either case the pre-existing value is totally superfluous if the collapse process is irreversible and indeterminist.
The notion that probabilities in QM result from the uncertainty principle has always been a minority view it appears, as most authors do not mention anything like it (this question has a mode of 00 in all samples). There is still some support for the notion that the uncertainty principle is a consequence of measurement disturbance, with assent frequency $>0.71$ outside the ‘modern’ sample where it is only $0.48$. This indicates that affirmation of this proposition may be falling out of the orthodoxy. Interestingly, the sample with most affirmation for this last proposition is the “Wikipedia notable” one, with $0.79$ assent ($p=1.17\times 10^{-3}$).
Thus, for the orthodoxy, we do not reveal the values of studied quantum systems but force them to assume definite values (or alter them) via some uncontrolled disturbance attached to the abstract notion of measurement. The “uncontrolled” aspect of this is vital to square the statistical nature of QM with the proposition that quantum states apply to individual instances of systems. Here we see emergence of anti-realism as a pillar of the orthodoxy. This is entailed by the fact that we are not exploring the pre-existing properties of quantum systems so much as forcing them into altered states under measurement.
To better understand the development of the orthodoxy over time we can also consult the mean responses in each sub-sample, as in Figure 2. Here we can see the modern sample has a consistently lower mean response to the vast majority of propositions. The most significant difference occurs for the “single particle uncertainty” proposition, with a $40$% reduction in the mean value relative to the other samples. This reduction is twice the difference in assent frequency between the samples, indicating the presence of active dissent from the proposition. This cannot be seen from the assent frequency, as $-1$ and 00 have the same effect on the frequency of $1$. The proposition of “ontological completeness” sees a decrease in mean response of around 20% in the ‘modern’ sub-sample that is not evident in response frequency, again showing evidence of reduced consensus on this point. The “uncertainty from disturbance” proposition also has a mean response decline of $30$%. However, this correlates well with the assent frequency, indicating only that this idea is increasingly not mentioned in more recent works. These demonstrate that there are propositions that were previously orthodox (with assent frequencies above $0.7$ in other samples) that have become more controversial over time. Some less popular positions (“probability from uncertainty” and “no trajectories”) also exhibit evidence of increased controversy in the modern sample, showing smaller mean values than even other samples where the mode is similarly 00.
Figure 2: Average values of the responses to each question. The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), and modern (grey).
Finally, we have Figure 3. This displays the “orthodoxity” of each book, i.e. how close they are to the averaged responses over the full sample. This is determined following
$O=\frac{\sum_{i}^{N}w_{i}a_{i}}{\sum_{i}^{N}w_{i}}\;,$ (1)
where the sum runs over the propositions, $w_{i}$ is the average response value, and $a_{i}$ is that of the reference work in question. Note that we exclude propositions that do not have assent frequency $>0.6$.
The average orthodoxity is $74\pm 25$%, indicating a robust degree of agreement on average. For convenience we order Fig. 3 by year of first publication. This allows us to determine if there are any trends in the orthodoxy. What we see is quite remarkable. Early works adhere very closely to the orthodoxy (as might be expected). The period around the early 90’s shows a fluctuation of far lower agreement with the orthodoxy. However, the orthodoxy reasserts itself after this point (but slightly more weakly than in the earlier works). The 90’s dip in the orthodoxity is constituted by the works: Peres [29], Moore [27], Griffiths [28], Mandl [30], Rae [13], and Sakurai [33]. Overall, there is some evidence for an evolutionary trend in the orthodoxy, but it must be admitted that modern works quite closely follow the interpretation put forward by their predecessors. Despite this as an overall statement, we can see recent dissent from Weinberg [17] and some clearly unorthodox positions in both Binney & Skinner [18] and Griffiths [28]. This is of interest given the prominence of the first’s author and that the latter two also appear on the “Wikipedia Notable” list. It is also interesting that the modern reduction in orthodoxity (even if unspectacular on average) revolves around the questions of the wave-like behaviour of individual particles. Something that is denied by Weinberg, Binney & Skinner, Griffiths, and Sakurai [17, 18, 28, 33], these being the works with some of the lowest orthodoxity. The lack of orthodoxity for Peres [29] is rooted in the author’s instrumentalism (being far closer to the version of a Copenhagen viewpoint as understood by Stapp [2]).
For additional works that were not included in the sample we find the following orthodoxity measures: Ballentine [12] $-0.89$, Home [4] 0.44, Hannabuss [14] 0.43, and Nielsen & Chuang [15] 0.44. The last of these, Nielsen & Chuang [15], only shows a lower orthodoxity by virtue of never mentioning the topic of waves. Their strong endorsement of Peres [29] and Sakurai [33] might indicate little sympathy with this aspect of the orthodoxy, but the clarity of this is marred by also endorsing Cohen-Tannoudji [36]. Interestingly, Home’s characterisation of the orthodoxy [4] appears incomplete at $0.44$. Additionally, this is only when one concentrates on the aspects attributed to the non-ensemble case. Home considers ensemble views as part of the orthodoxy on the basis that they view the formalism as being complete (there are no supplemental entities needed to explain what is going on). This principle as the central pillar of the orthodoxy is apriori dubious. This is because the notion of collapse is a supplement to the basic mathematical formalism [49, 12, 17], in the form of an additional physical/quasi-physical process with attendant projection postulate. Since Copenhagen-like interpretations hinge so heavily on collapse, their popularity implies that Home’s primary condition of orthodoxy becomes more doubtful. Interestingly, Ballentine [12], who would express an orthodox view according to Home, achieves an orthodoxity of $-0.89$ here, dissenting from (or not addressing) every proposition considered. This is because Home’s rationale for including ensemble views in the orthodoxy may not be well justified. Home argues that, to enjoy a complete formalism, the ensemble must be supplemented by hidden variables that explain the assignment of definite values to individual realisations. However, an ensemble view only requires statistical completeness, $\psi$ can answer all statistical questions, not true ontological completeness. In this regard Home forces ensemble views to be able to determine the behaviour of single instances, in contradiction with their basic premise. In addition, no such definite assignment is required of other statistical theories in science (which usually exist as a limit of some other, more detailed theory). We do not require statistical mechanics to explain how velocities are assigned to the molecules in a gas, for instance.
Figure 3: The “orthodoxity” or degree to which each studied source agrees with the averaged responses to each proposition. This is ordered by date of first publication.
We can now ask how far the revealed orthodoxy diverges from the idea that it can be defined entirely by the ontological primacy of the quantum state (as advocated by Home [4]). The first element that brings this into question is the “non-ensemble $\psi$” proposition. This is a vital part of the orthodoxy (stronger even than ontological completeness) and it does not follow from $\psi$ being ontologically complete. There being no reason to preclude a fundamentally indeterminist viewpoint which does not require single realisations to be associated to $\psi$. If this is the case then there are a cascade of propositions that must follow, as they all depend on “non-ensemble $\psi$” for their relevance. These are “single particles are wave-like”, “states collapse on measurement”, “measurement inherently disturbs the system”, and “measurement creates observed values”. This indicates that a single proposition reduction of the orthodoxy is unviable. The two main pillars that are empirically revealed are “measurement disturbs systems” and “states do not just describe ensembles”. Notably, the significance of consensus around ontological completeness has declined for modern works, as a result of some authors [24, 27] instead positing “statistical completeness” or openly suggesting a more general theory should be found [17].
6 Conclusion
When we consider a sample of $33$ textbooks, aimed at both graduate and undergraduate students of quantum mechanics, we find a robust orthodoxy emerging. This is powerfully illustrated by the average orthodoxity of the sources being $74$% with a standard deviation of only $25$%. Given that the studied sample spans nearly a century, this provides good evidence for a mostly unchanged consensus on the interpretation of quantum mechanics. In modern works there is slightly less support for the notion that individual particles are wave-like, but large reductions in support for the notions that the uncertainty principle applies to individual particles, and that quantum uncertainty results from measurement disturbance.
The revealed orthodoxy strongly resembles many aspects of Copenhagen-like frameworks. Particularly, there is powerful agreement (assent frequency $0.92$) that measurement inherently disturbs quantum systems, single instances of systems are individually described by a quantum state, rather than as an ensemble (assent frequency $0.95$), state collapse happens (assent frequency $0.77$) , that individual particles are wave-like (assent frequency $0.85$), and that quantum states are ontologically complete (assent frequency $0.85$). These positions are further supplemented by the existence of a wave-particle duality (assent frequency $0.79$) and that measurement actually creates or alters the definite property values being measured (assent frequency $0.71$). More contentious, but still popular, are the notions that the uncertainty principle arises from measurement disturbance (assent frequency $0.71$), and that individual particle measurements are subject to the uncertainty principle (assent frequency $0.82$). These assent frequencies have a very small probability to arise by chance when drawing samples from a random binomial distribution, indicating a level of statistical robustness in the observed consensus.
It is clear that the orthodoxy favours the view of the quantum state as a complete description of the actual state of affairs, rather than the more instrumental view attributed to Copenhagen by authors like Stapp [2]. In this regard, the views of the orthodoxy espoused by Home [4] and Jammer [3] are vindicated. The work most embodying this kind of orthodoxy would be Peres [29], with similarities in Mandl [30], Nielsen & Chuang [15], Hannabuss [14], and Sakurai [33]. However, it is not clear that all the logical consequences of the single proposition view are believed by the physics community (and it is clear there are wave-related addenda). Thus, the orthodoxy’s characterisation requires more nuance than being distilled to a single point about the status of quantum states. It is clear that, examining all the propositions in this work, there is no room in the orthodoxy for ensemble views, entirely in opposition to the position of Home. In fact, it is a strong requirement of the orthodoxy that quantum states describe single instances, rather than ensembles.
To an orthodox quantum mechanic, individual particles exist in wave-like superpositions, when left to their own devices, and assume their observed properties (and appear as particles) when forced to by outside interference. When this happens, the individual particle’s properties are subject to the uncertainty principle as a form of measurement uncertainty and the state is inevitably changed by the uncontrolled disturbance inherent in measurement. Perhaps understandably, few textbook authors state the anti-realist aspects of these views with complete openness. For instance the “measurements create observed values” proposition is explicitly endorsed by only around 50% of the texts but remains a plausible corollary of positions taken by another 40%.

### v2

@@
 However, we adopt a different approach to that normally used.
-Rather than carefully surveying the precise details of the thoughts of Bohr and Heisenberg, we extract an orthodoxy empirically.
-To do this we review a collection of $33$ textbooks on quantum mechanics, encompassing the most popular and prominent works of this nature.
+Rather than surveying physicists, or poring over the precise details of the thoughts of Bohr and Heisenberg, we review a collection of 42 textbooks on quantum mechanics, encompassing the most popular and prominent works of this nature.
 We then gauge their response to 12 propositions to build up a picture of exactly what is believed by an orthodox quantum mechanic.
-We demonstrate that this orthodoxy is largely unchanged over the past century, with some interesting emerging deviations, and has many aspects of Copenhagen-like viewpoints.
-However, it is more nuanced than some reductive characterisations that condense it down to the ontological primacy of the quantum state.
-The revealed orthodoxy has two main pillars: measurement inherently disturbs quantum states and these states refer to individual instances, not ensembles.
-More fully it entails that individual particles exist in wave-like super-positions and present particle behaviours only when forced to by outside influences.
-The act of measuring such a system inherently changes its state in a random fashion, manifesting in a form of measurement error that corresponds to the uncertainty principle.
+We demonstrate that this orthodoxy has many aspects of Copenhagen-like viewpoints, but also shows some interesting emerging deviations.
+Moreover, it is more nuanced than some reductive characterisations that condense the orthodoxy down to the ontological primacy of the quantum state.
+The revealed orthodoxy has two consistent pillars: measurement inherently disturbs quantum states and these states refer to individual instances, not ensembles.
+More fully, it entails that individual particles exist in wave-like super-positions and present particle behaviours only when forced to by outside influences.
+The act of measuring such a system inherently induces random changes in its state, manifesting as a form of measurement error that corresponds to the uncertainty principle.
 This implies that measurement does not reveal underlying values of quantum properties.
 keywords:
-Quantum foundations; Quantum interpretation
+quantum mechanics, quantum interpretation, quantum foundations
 1 Introduction
 How do scientists understand quantum mechanics?
-It is commonly asserted that the most popular interpretation of quantum mechanics (QM) is the “Copenhagen interpretation”.
-This is, however, notoriously vague [1, 2, 3, 4] and its characterisation tends to descend into hyperfine readings of Bohr and Heisenberg [2].
-Rather than characterise this Copenhagen model, it seems more productive to instead determine what the consensus within the physics community actually is.
-In other words, we can look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
-The question of what the orthodox view of quantum theory is has, of course, received previous attention in the literature.
+The most popular interpretation of quantum mechanics (QM) is usually found in surveys to be “Copenhagen interpretation” [1, 2].
+However, the definition of this viewpoint is notoriously vague [3, 4, 5, 6], its characterisation tending to descend into hyperfine readings of Bohr and Heisenberg [4].
+Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [5].
+These include the completeness of the state and the projection postulate.
+To produce something like a Copenhagen interpretation, this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [7] described as “doing epistemology without ontology”.
+This all results in a similar view to Home, that it is the primacy of the quantum state that typifies the orthodox view.
+Interestingly, Stapp argues [4] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities.
+This seems somewhat difficult to square with the notion that the state is the most complete specification of the real situation, without some explanation of why additional information is inaccessible.
 Home [6] argues that the primary characteristic of an orthodox viewpoint is that the state is framed as a complete description of a quantum system.
@@
 Home refers to Feyeraband’s encapsulation of the standard position as being that quantum systems possess no definite properties over and above those that can be derived from the state itself [8].
-Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [3].
-These include the completeness of the state and the projection postulate.
-To produce something like a Copenhagen interpretation this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [6] described as “doing epistemology without ontology”.
-This all results in a similar view to Home, that it is the primacy of the quantum state that typifies the orthodox view.
-Interestingly, Stapp argues [2] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities.
-This seems somewhat difficult to square with the notion that the state is the most complete specification of the real situation, without some explanation of why additional information is inaccessible.
 These points reveal a tension in descriptions of a quantum orthodoxy: a Copenhagen-like view wants to claim a “completeness” or “finality” for QM (the “end of the road hypothesis” as Popper terms it [9]).
@@
 It seems clear that a follower of an allegedly Copenhagen viewpoint could pick and choose between these two, somewhat opposed, strands of thought.
-Clearly, it is not sufficient to characterise the orthodoxy of quantum interpretation by the single question “is the quantum state a complete description of reality?” At least not without presupposing the logical coherency of the orthodoxy and erasing the nuance in what we mean by “complete description”.
-This is typified by Home’s collapse of ensemble views into the orthodoxy.
-We must note that there is, so far, no experiment whose statistics are not accurately predicted by quantum mechanics.
-However, this does not itself entail all of the views associated with Bohr and Heisenberg that both Home and Jammer link to a standard viewpoint.
-What we need is a more detailed breakdown of interpretative beliefs that constitute the orthodoxy, divorced carefully from the common formalism.
-How then do we study the beliefs of physicists as a community?
-The answer is that we turn to textbooks.
+Rather than characterise this Copenhagen model, it seems more productive to instead determine what the consensus within the physics community actually is.
+With a particular focus on answers to questions about quantum phenomena, along the lines of those used in [1].
+Existing surveys of physicists tend to reveal a very fragmented community [1, 2], instead we will look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
+The target of our survey will be textbooks.
 The reason for this is that these works are presented for the education of a new generation of scientists.
@@
 Thus, even the beliefs of established researchers in quantum interpretation may be strongly correlated with the orthodoxy found in textbooks.
-It is therefore suggested that one cannot completely dismiss the interpretive remarks made in reference works on this subject.
-They fulfill a vital function in passing along a common understanding of quantum mechanics that is insufficiently conveyed by the mathematical formalism alone.
-This work attempts an empirical investigation of the quantum orthodoxy via a survey of popular and widely used textbooks on quantum mechanics, for both graduates and undergraduates.
-This encompasses $33$ works, with dates of first publication spanning from 1930 to 2019.
-The sample starts with a set of works listed as “notable” on Wikipedia [^1] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics and expands it via textbooks that rank among Amazon best-sellers or are written by notable authors.
+It is therefore suggested that one cannot dismiss the interpretive remarks made in reference works on this subject.
+They fulfil a vital function in passing along a common understanding of quantum mechanics that is insufficiently conveyed by the mathematical formalism alone.
+This work attempts an empirical, if unavoidably subjective, investigation of the quantum orthodoxy via a survey of popular and widely used textbooks on quantum mechanics, for both graduates and undergraduates.
+Since we are investigating interpretation (or perhaps beliefs) this cannot be wholly objective.
+First, one must select which books to survey, there being no means of defining a representative sample.
+Our sample encompasses $42$ works, with dates of first publication spanning 1930 to 2025.
+The sample starts with a set of works listed as “notable” on Wikipedia [^1] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics and expands it via textbooks that rank among Amazon bestsellers or are written by notable authors.
+This is not a rigorously chosen sample, if such a thing could even be defined in this context.
+However, we are searching for points of consensus, so a large sample containing the most popular works should be sufficient for our purposes.
+Additionally, bias from sample selection is likely to be less of an issue here than in surveys of physicists themselves.
 To determine an orthodoxy empirically, we require some kind of framework of positions that might be part of the orthodoxy.
-Thus, we establish a set of 12 propositions, divided into three categories.
-The first of these encompasses how we understand the state in quantum mechanics.
-Here we identify three propositions that provide a Copenhagen-like framework: “the state is ontologically complete”, “the state does not merely describe an ensemble of identically prepared systems”, and “the state collapses upon measurement”.
+Once again, this injects subjectivity into our analysis, as it is not clear how we might choose these positions.
+We settle on a set of 12 propositions, divided into three categories.
+These are selected to at least cover the major assertions of a Copenhagen-like framework and show strong overlap with surveys [1].
+The first category explores how we understand the quantum state, with the propositions: “the state is ontologically complete”, “the state does not merely describe an ensemble of identically prepared systems”, and “the state collapses upon measurement”.
 The second category revolves around the nature of particles with the propositions: “individual particles have wave-like properties”, “particle-like entities are described by a wave-particle duality”, “particles cannot simultaneously possess definite values of non-commuting observables”, “particles do not follow classical-like trajectories”, and “there is no adequate single picture of particle or wave”.
@@
 These will all be elaborated on in later sections of this paper.
-Each work is scrutinised and a trinary value is assigned to their agreement with each proposition (an in-between value is used when agreement can be strongly inferred from the context rather than explicitly stated).
-It can then be determined to what degree the works assent to each proposition on average and, by analysing the statistics, a set of propositions that robustly represent a consensus view point can be extracted (should one exist).
+Each work is scrutinised, and a trinary value is assigned to their agreement with each proposition (an in-between value is used when agreement can be strongly inferred for consistency with other positions taken).
+It can then be determined to what degree the works assent to each proposition on average and, by analysing the statistics, a set of propositions that represent a consensus viewpoint can be extracted (should one exist).
 This paper is structured as follows: we first lay out the candidate propositions in section 2, then we detail the sample of textbooks considered in section 3, display example analyses of texts in section 4, and discuss the statistical properties in section 5.
@@
 Within each category we include propositions that source works can either agree with or dissent from.
-The selection of these questions was designed partly apriori and partly by assertions encountered in some of the works.
-We can then see why there is no mention of most of the vast multitude of interpretative approaches: they tend to feature only within scientific papers, rather than in textbooks designed to educate new generations of scientists.
+The selection of these questions was designed partly apriori and partly by assertions encountered in some works.
 We reiterate that the choice of textbooks over research articles is made because we feel that the most distilled consensus is to be found in textbooks, which typically are formulated around what is considered “established knowledge” within a given field.
 It is entirely possible that this question set is incomplete (or contains some redundancy).
-However, we will see that the results deduced are quite statistically robust, even with the small sample size.
+We also note that our choices cover questions asked of working physicists in surveys [1, 2].
+Notably, the aforementioned [1] also includes “are values are defined prior to measurement?” (“measurement creates values”), “is randomness inherent in QM?” (strong overlap with “ontological completeness”), “where is the electron in a hydrogen orbital?” (“no trajectories”), “is macroscopic superposition possible?” (linked to “the state doesn’t just describe an ensemble”), and a question on the solution of the measurement problem is covered by the entire set of questions on why QM is statistical.
+However, we do not include questions about the role of the observer or the implications of Bell inequality violations.
+These latter points are highly contentious in the literature and do not receive consistent coverage in textbooks.
+The cases of incomplete overlap are down to the luxury of expositive works like textbooks.
+In that, our questions can be simpler, more numerous, and less nuanced than those in [1, 2], as they need not be processed in real time by the respondent.
 2.1 What is the quantum state?
@@
 This stands opposed to cases like the de Broglie-Bohm interpretation, where the state is supplemented by hidden variables and a pilot wave equation.
-An ensemble interpretation might deny this proposition, but affirm that QM is statistically complete, not because it supplements the formalism or metaphysics, rather it leaves open that QM may be a quantum analogue of statistical mechanics.
 Another viewpoint that would disagree with this proposition is one that assigns only epistemic value to the state.
-For the purposes of this work, we will take a response such as “the state represents all we can know about a system” as affirming the proposition, unless they supply some explanation of why the full ontology is not accessible.
+On the other hand, an ensemble interpretation might deny this proposition, but affirm that QM is statistically complete, not because it supplements the formalism or metaphysics, rather it leaves open that QM may be a quantum analogue of statistical mechanics.
+A question arises here in how we delineate epistemic viewpoints.
+In this work we take “the state contains all that can be known about the system” as assent to ontological completeness.
+The “known” aspect might make this seem questionable.
+However, we focus on the “can be known”, which implies an ontological limit on knowledge.
+If a work asserts that the state is a “calculational tool only” or that it “only characterises our knowledge of the system”, we will deem this dissent from ontological completeness.
 2.1.2 The state does not just describe ensembles
-Here we mean that $\psi$ describes single realisations of a given system (shorthand “single instance $\psi$”).
+Here we mean that $\psi$ describes single realisations of a given system (shorthand “non-ensemble $\psi$”).
 It is not limited to the statistics of an ensemble of similarly prepared systems.
@@
 2.2.2 Wave-particle duality is necessary
-This is an assertion of the fundamental importance of wave-like behaviour, even if the wave-like behaviour is not displayed by individual particles.
-Unobserved particle-like entities behave like waves, or have wave-like probability densities, but like particles when we detect them.
-Thus, we have dual, mutually exclusive, descriptions of the same experiments.
+This is an assertion of the fundamental importance of wave-like behaviour, even if individual particles do not display wave behaviour.
+We have dual, mutually exclusive, descriptions of the same experiments.
 This can be viewed as a weaker form of the preceding proposition.
@@
 That is, if I measure the position of a particle I cannot also know its momentum simultaneously.
-More generally: systems cannot posses definite values of non-commuting observables simultaneously.
-This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble [7].
+More generally: systems cannot possess definite values of non-commuting observables simultaneously.
+This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble of measurements performed on similarly prepared systems [9].
 2.3 Why are our predictions statistical?
-When we conduct quantum mechanical experiments we require measuring over an ensemble in order to compare to the probabilistic predictions of the theory.
-If our interpretation of the quantum state is not ensemble-based, we have to explain why this form of measurement should be the status quo.
+When we conduct quantum mechanical experiments we require measuring over an ensemble of similarly prepared systems in order to compare to the probabilistic predictions of the theory.
+If our interpretation of the quantum state is not ensemble-based or epistemic, we have to explain why this form of measurement should be the status quo.
 These propositions examine some explanations (and related ideas) that arise in the studied works, in order to determine which ones may belong to the orthodoxy.
@@
 Measurement inherently changes the state of a system while measuring it (except in the case of eigenstates).
-A viewpoint that denies this would note that non-interactive measurements exist in QM or that measurement has no general abstract definition.
+A viewpoint that denies this would note that non-interactive measurements exist in QM or that measurement has no abstract definition.
 2.3.2 Uncertainty results from disturbance
@@
 This results in a random statistical scatter in our measurements over an ensemble of similarly-prepared systems.
+An opposing view would be that uncertainty is present in quantum states even during Schrödinger evolution, so it cannot be associated with disturbance, or that the uncertainty principle can be validated on separate, similarly prepared samples (one set for $\hat{x}$ and the other for $\hat{p}$ for instance).
 2.3.3 Uncertainty leads to probabilities
@@
 The consequences of this is that measurements have probabilistic results, even though quantum mechanics is not formulated as fundamentally statistical.
-One might deny this if one recognises that the uncertainty principle is a consequence of quantum rules, rather than a postulate, or if one correctly notes the uncertainty principle is formulated in terms of statistical observables only.
+One might deny this if one recognises that the uncertainty principle is a consequence of quantum rules, rather than a postulate, or if one notes the uncertainty principle is formulated in terms of statistical observables only.
 2.3.4 Measurement creates observed values
-The observed values of quantum properties are the result of an interaction between apparatus and system, in the sense that they do not reflect underlying pre-measurement values in cases that are not eigenstates of observable being measured.
+The observed values of quantum properties are the result of an interaction between apparatus and system that alters the measured property.
+In this sense, they do not reflect underlying pre-measurement values in cases that are not eigenstates of the observable being measured.
 This stands in opposition to the usual scientific idea of measurement: that we are generally discovering the values of pre-existing properties.
-This assertion can take two forms: a strong form, where it is asserted that quantum systems do not actually possess definite properties till measured, and a weaker form.
-In the latter, authors might take the position that measurement merely changes pre-existing values.
-Both of these result in the pre-measurement value being indeterminate and irrelevant.
+This assertion can take two forms: a strong form, where it is asserted that quantum systems do not actually possess definite properties till measured, and a weaker form where measurement merely changes pre-existing values.
+Both of these result in the pre-measurement value being indeterminate and potentially irrelevant.
 3 Sample of studied works
-To perform this study we examine the $33$ popular textbooks listed in Table 1.
-These range in date of first publication from 1930 to 2019.
+To perform this study we examine the $42$ popular textbooks listed in Table 1.
+These range in date of first publication from 1930 to 2025.
 This large time span is of utility, as it allows us to determine whether there is any evolution within the consensus view of quantum mechanics.
-Additionally, we consider subdivision of these works into graduate ($N=14$) and undergraduate ($N=19$), as determined from the statements of the target audience in the work itself.
-These works were chosen to encompass the Wikipedia list on ‘Notable quantum mechanics textbooks’ [^2] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics ($N=17$) as well as to include as large a sample as feasible.
-To detect variation over time we also consider a sub-sample of works first published after 1980 ($N=24$).
-Note that Omnés work is included, even though it is a textbook that advocates against the orthodoxy in favour of “consistent histories”, as it provides its own description of the orthodox positions.
-In the case of [13], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the ‘conventional” approach.
+Additionally, we consider subdivision of these works into graduate ($N=17$) and undergraduate ($N=25$), as determined from the statements of the target audience in the work itself.
+These works were chosen to encompass the Wikipedia list on ‘Notable quantum mechanics textbooks’ [^2] https://en.wikipedia.org/wiki/List_of_textbooks_on_classical_mechanics_and_quantum_mechanics ($N=18$) as well as to include as large a sample as feasible.
+To detect variation over time we also consider three subsamples: works pre-1980 ($N=15$), post-1980 ($N=27$), and post-2000 ($N=17$).
+Note that Omnés work is included, even though it is a textbook that advocates against the orthodoxy in favour of “consistent histories”, as it provides its own description of the orthodox positions which we use for our study.
+In the case of [15], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the “conventional” approach.
+Note that page references for [16] are in reference to the online lecture notes on which it is based [^3] https://www.damtp.cam.ac.uk/user/tong/books/quantum.html, due to the recency of the text.
 We will also make reference to a set of 4 works whom we do not include in any of the samples: [6, 17, 18, 14].
-The first is excluded as it is a discussion of interpretations but will provide a useful point of comparison.
-The second two are analysed out of curiosity but are not targeted squarely as part of a physics syllabus.
+The first is excluded as it is a discussion of interpretations, but will provide a useful point of comparison.
+The second two are analysed out of curiosity, but are not targeted squarely as part of a physics syllabus.
 The final one will be analysed and found to be entirely outside the orthodoxy on almost every point of contention (so it is treated as an outlier).
@@
 Title Author Year Ref Level Wiki
+Lecture notes on quantum mechanics D. Tong 2025 [16] U N
 Basic quantum mechanics K. Tamvakis 2019 [19] U N
+A first introduction to quantum physics P. Kok 2018 [20] U N
 Lectures on quantum mechanics S. Weinberg 2013 [21] G N
@@
 Quantum mechanics D. McIntyre et al 2012 [24] U Y
+Quantum mechanics: theory and experiment M. Beck 2012 [25] U N
 Quantum mechanics: a new introduction K. Konishi & G. Paffuti 2009 [26] U N
+Quantum mechanics for scientists and engineers D. Miller 2008 [27] U N
+Quantum mechanics: a textbook for undergraduates M. Jain 2007 [28] U N
 Introduction to quantum mechanics H. Müller-Kirsten 2006 [29] U Y
 Quantum physics M. Le Bellac 2006 [30] G N
+Lectures on quantum mechanics J. Basdevant 2005 [31] G N
 Introduction to quantum mechanics A. Philips 2003 [32] U N
 Quantum Mechanics: Concepts and Applications N. Zettili 2001 [33] U Y
+Elements of quantum mechanics M. Fayer 2001 [34] U N
 Understanding quantum mechanics R. Omnés 1999 [35] U N
@@
 Quantum mechanics C. Cohen-Tannoudji 1977 [45] G Y
-Quantum physics of atoms, molecules, R. Eisberg & R. Resnick 1974 [37] U Y
-solids, nuclei and particles
+Quantum physics of atoms, molecules, solids, nuclei and particles R. Eisberg & R. Resnick 1974 [46] U Y
 Quantum physics S. Gasiorowicz 1974 [47] U Y
 Feynman lectures on physics volume 3 R. Feynman 1965 [48] U Y
+Quantum mechanics A. Davydov 1965 [49] G Y
 Elementary quantum mechanics P. Fong 1962 [50] U N
+Quantum mechanics E. Merzbacher 1961 [51] G N
 Introduction to quantum mechanics R. Dicke & J. Wittke 1960 [52] G N
@@
 These are chosen because they all share a common case of having to infer agreement with at least one proposition (usually “measurement creates observed values”).
-This will allow the reader to determine that we are carefully evidencing degrees of assent or dissent in each case.
+This allows the reader to determine that we are carefully evidencing degrees of assent or dissent in each case.
 Note that the sections below leave out propositions that are unaddressed by the given text.
@@
 For “uncertainty linked to disturbance” we find evidence for endorsement on page 3 (“We rarely know the exact state that our measuring instrument is in before we bring it into contact with the system we have measured, so the result of the measurement of the atom would be uncertain even if we knew the precise state that the atom was in before we measured it, which of course we do not”), page 28 (a discussion of the uncertainty principle in terms of measurement uncertainty) which is subsequently reinforced on pages 130 and 131.
-For single particle uncertainty we find on page 28 “the position of a particle that has well defined momentum is maximally uncertain” as evidence for adherence to this position.
+For single particle uncertainty we find on page 28 “the position of a particle that has well-defined momentum is maximally uncertain” as evidence for adherence to this position.
 For “particles are wave-like” we find a discussion opposing this on page 7, particularly “we cannot consistently infer from this correspondence that particles are manifestations of waves”.
-For “collapse”, the work opposes this idea explicitly stating collapse is not physical on page 15 and that the hypothesis is false on page 130.
+For “collapse”, the work opposes this idea, explicitly stating collapse is not physical on page 15 and that the hypothesis is false on page 130.
 For “particle-wave duality” the authors never mention the concept explicitly but already oppose “single particles are wave-like” outside of quantum field theory contexts (page 7).
-For “non-ensemble $\psi$” we find on page 14 “in quantum mechanics an individual photon has an amplitude to be linearly polarised in a any chosen direction and an amplitude to be circularly polarised in a given sense”, so states correspond to individual system realisations.
+For “non-ensemble $\psi$” we find on page 14 “in quantum mechanics an individual photon has an amplitude to be linearly polarised in any chosen direction and an amplitude to be circularly polarised in a given sense”, so states correspond to individual system realisations.
 This is reinforced on page 71 where they state “In quantum mechanics, complete information about any system is contained in its ket $\left|\psi\right$>”.
@@
 On page 5 uncertainty is described in terms of single particle measurement precision or “if we make a measurement on any object and determine its position along a given direction with uncertainty $\Delta x$, its momentum along this direction will necessarily be known with uncertainty no less than $\frac{h}{4\pi\Delta x}$”, an indication of “single particle uncertainty”.
-On page 6 the de Broglie relation is discussed as the wave aspect of a single particle, implying individual particles are wave-like.
+On page 6, the de Broglie relation is discussed as the wave aspect of a single particle, implying individual particles are wave-like.
 Pages 86 and 87 discuss a non-unitary evolution during measurement and call this collapse, this re-iterated on page 442.
@@
 Page 5 describes the uncertainty principle in terms of knowledge of $x$ and $p_{x}$ “thus, viewed from a classical viewpoint, in a microscopic system, complete knowledge of $x$ and $p_{x}$ is impossible”, we infer this is indicating that probabilities result from incomplete knowledge implied by the uncertainty principle.
-For “measurement observed creates values” we note that it is stated on page 86 that a superposition represents the pre-measurement state, since this cannot have a definite property value we must conclude that definite property values are created, or at least changed, during measurement.
-This will satisfy an assent to “measurement observed creates values” as a pre-measurement state with definite values cannot be recovered as the author describes measurement as “acausal” on page 87.
+For “measurement creates observed values”, we note that it is stated on page 86 that a superposition represents the pre-measurement state, since this cannot have a definite property value we must conclude that definite property values are created, or at least changed, during measurement.
+This will satisfy an assent to “measurement observed creates values” as a pre-measurement state with definite values cannot be recovered, as the author describes measurement as “acausal” on page 87.
 4.3 Zettili
 Next we consider Quantum Mechanics: Concepts and Applications by Zettili [33].
-On page 24 “measurement interferes with the state of microscopic objects” and “it is the effects of interference by the equipment on the system which is the essence of quantum mechanics” on page 165 demonstrate support for measurement disturbance.
+On page 24, “measurement interferes with the state of microscopic objects” and “it is the effects of interference by the equipment on the system which is the essence of quantum mechanics” on page 165 demonstrate support for measurement disturbance.
 On page 169 uncertainty is described as a result of measurement “perturbing” the system.
@@
 On page 158 collapse is included in the postulates of quantum mechanics.
-On page 26 “Microscopic systems, therefore, are neither pure particle nor pure waves, they are both” as well as direct invocations on the same page, suffices for duality and complementarity.
+On page 26 “Microscopic systems, therefore, are neither pure particle nor pure waves, they are both”, as well as direct invocations on the same page, suffices for duality and complementarity.
 For ontological completeness, page 158 asserts “the state contains all the needed information about the system”.
@@
 4.4 McIntyre, Manogue, & Tate
-We now turn to Quantum mechanics by McIntyre et al [20].
+We now turn to Quantum mechanics by McIntyre, Manogue, and Tate [24].
 On page 8 we find that measurements disturbing systems is a “key feature” of quantum mechanics.
@@
 A confusing discussion on 207 about the double-slit seems to suggest both that a wave picture is just for finding particle probabilities but also that the waves have to actually interfere (passing through two slits).
-The same discussion refers to Feynman and Cohen-Tannoudji which seems to indicate the ultimate outcome is individual particles are wave-like too.
+The same discussion refers the reader to Feynman and Cohen-Tannoudji which seems to indicate the ultimate outcome is individual particles are wave-like too.
 Page 192 indicates we cannot know a particles trajectory but does not deny its existence.
@@
 Discussion of Schrödinger’s cat on 167 and 168 is illuminating in terms of the role of ensembles.
-Here it is asserted that only decoherence prevents superpositions at classical scales and interference is used to ascribe physical signficance to superposition, beyond mere statistics.
+Here it is asserted that only decoherence prevents superpositions at classical scales and interference is used to ascribe physical significance to superposition, beyond mere statistics.
 Thus, we see that the authors take superpositions to describe the state of individual realisations.
-Acceptance of real superpositions combined with the stated quantum postulates requires definite values are created by measurement.
+Acceptance of real superpositions combined with the stated quantum postulates requires observed values are created by measurement.
 4.5 Konishi & Paffuti
@@
 These indicate that quantum states do not apply to ensembles but to single realisations of systems.
-On page 29 the authors agree that probability derives from the uncertainty principle “the probabilistic nature of quantum mechanical pediction for the results of measurement is an inevitable consequence of Heisenberg’s uncertainty principle.” Although the discussion on 518 and 519 is ambiguous as to whether values are being revealed or created (the authors use “indefinite” properties of a superposition here to mean “unpredictable”), we must infer that they are at least being altered by measurement to comport with previous assertions.
+On page 29 the authors agree that probability derives from the uncertainty principle “the probabilistic nature of quantum mechanical prediction for the results of measurement is an inevitable consequence of Heisenberg’s uncertainty principle.” Although the discussion on 518 and 519 is ambiguous to whether values are being revealed or created (the authors use “indefinite” properties of a superposition here to mean “unpredictable”), we must infer that they are at least being altered by measurement to comport with previous assertions.
 On page 21 the authors state that “knowledge of the wave function amounts to complete knowledge of the state” which, combined with prior assertions fully supports ontological completeness.
@@
 Each work’s response to the propositions is characterised by a degree of agreement.
-This is kept as simple as possible with three principle values: $1$ indicates agreement, $-1$ dissent, and 00 as not mentioned.
-We expand this via either of ±$\pm 0.5$ to mean ‘inferred for logical consistency’.
+This is kept as simple as possible with three principle values: $1$ indicates agreement, 1-1 dissent, and 0 as not mentioned.
+We expand this via either of $\pm 0.5$ to mean ‘inferred for logical consistency’.
 Inference of agreement/dissent is only made if the case is strong (in order to allow for a conservative analysis).
 We proceed to study the responses by observing the frequency of agreement.
-Note that this computation counts ±$\pm 0.5$ responses as half a count when determining frequencies.
+Note that this computation counts $\pm 0.5$ responses as half a count when determining frequencies.
 Our raw data keeps the page references separate for each proposition (available on request).
@@
 […]
+Tong [16] 1 1 1 1 1 1 -1 0 1 1 0.5 1 2,3,4,5,13,14,16,17,18,21,78
 Tamvakis [19] 1 0 1 1 0 1 1 1 1 1 0.5 1 5,6,7,86,87,442
+Kok [20] 1 1 1 1 1 1 1 0 1 0 1 0 6,182,186,211,58,206,225,226
 Weinberg [21] 0 0 0.5 -1 0 -1 0 0 1 0 0 -1 21,25,82,83,88,95,97
@@
 Mcintyre et al [24] 1 0.5 1 1 -1 1 1 1 1 0 0.5 1 7,8,48,59,121,167,168,192,207,208
+Beck [25] 1 0 1 1 0 1 0 1 1 0 1 1 62,91,92,101,129
 Konishi & Paffuti [26] 1 1 1 1 1 1 1 0 1 1 0.5 1 6,21,23,25,26,28,29,518,519
+Miller [27] 1 0 1 1 0 1 1 0 1 0 1 1 6,9,74,83,84,425,450
+Jain [28] 1 1 1 1 0 1 1 1 1 0 1 1 61,77,83,86,89,209,210,214
 Müller-Kirsten [29] 1 0 1 1 0 0 1 1 1 1 1 1 9,13,17,19,20,169
 Le Bellac [30] 1 0 1 1 0 1 1 0 1 0 1 1 22,24,97,101,102,104,179,185
+Basdevant [31] 1 -1 1 1 1 1 1 1 1 0 1 1 3,17,20,35,43,47,73,79,171,420
 Phillips [32] 1 1 1 1 1 1 1 0 1 0 1 -1 6,11,13,16,21,27,50
 Zettili [33] 1 1 1 1 -0.5 1 1 1 1 0.5 0.5 1 18,24,25,26,28,30,51,158,165,169
+Fayer [34] 1 1 1 1 1 1 1 0 1 0 1 1 1,2,3,5,32,33,42,55
 Omnés [35] 1 1 1 1 1 1 1 1 1 1 1 1 31,45,47,49,53,91
@@
 Feynman [48] 1 1 1 1 1 1 1 1 1 1 1 1 1-(6,7,9,10,11),2-(1,2,3,8,10)
+Davydov [49] 1 0 1 1 1 1 0 1 1 0 1 1 2,3,4,6,33,37,40,51
 Fong [50] 1 1 1 1 0 1 1 1 1 1 1 1 28,32,50,51,52,53,114,116,123,124
+Merzbacher [51] 0 0 1 1 1 -1 1 1 1 0 1 1 3,4,6,9,17,18,28,408
 Dicke & Wittke [52] 1 1 1 1 0 1 1 1 1 0 1 1 14,15,16,24,27,40,74,91,110,115,120
@@
 In Figure 1, we see the agreement frequencies for each of the questions in turn.
-We display these in terms of 4 samples: the full sample (black), undergraduate books (yellow), graduate texts (blue), works on the ‘Notable books’ list from Wikipedia (red), and ‘modern’ works (gray).
-Full information on the degree of agreement, and the $p$ value computed from a binomial distribution are included in Tables 3 to 7.
-These $p$ values are drawn from the binomial distribution and represent the probability of the given degree of assent to arise by chance given the size of the sample.
-We choose a random binomial distribution because we consider a zero consensus result would be equal frequencies of assent ($1$) and non-assent (where we group 00 and $-1$ together).
-The grouping is done because no assertion in the sample has a consensus answer of 00 or $-1$.
-Thus, we take a high $1$ frequency as agreement with this assertion is “in the orthodoxy”.
-We also include a $\sigma$ value, the number of standard deviations of the binomial distribution that the assent frequency corresponds to.
-Figure 1: Mode frequency of the responses to each question. The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), and modern (grey).
-Table 3: Degree of consensus from the full sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
-Assertion Affirmation $p$ $\sigma$
-Collapse 0.77 1.62e-04 3.1
-Non-ensemble psi 0.95 1.16e-10 5.2
-Ontologically complete 0.85 5.46e-06 4.0
-No trajectories 0.45 3.64e-01 0.5
-Complementarity 0.53 2.43e-01 0.3
-Wave-particle duality 0.79 1.62e-04 3.3
-Single particle uncertainty 0.82 3.31e-05 3.7
-Single particle wave-like 0.85 5.46e-06 4.0
-Measurement creates values 0.71 2.28e-03 2.4
-Probability from uncertainty 0.35 8.14e-02 1.7
-Measurement disturbs system 0.92 6.54e-08 4.9
-Uncertainty linked to disturbance 0.71 2.28e-03 2.4
-Table 4: Degree of consensus from the undergraduate sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
-Assertion Affirmation $p$ $\sigma$
-Collapse 0.76 9.61e-03 2.3
-Non-ensemble psi 0.97 1.91e-06 4.1
-Ontologically complete 0.79 2.21e-03 2.5
-No trajectories 0.39 3.24e-01 0.9
-Complementarity 0.55 1.80e-01 0.5
-Wave-particle duality 0.84 3.64e-04 3.0
-Single particle uncertainty 0.82 3.64e-04 2.8
-Single particle wave-like 0.89 3.81e-05 3.4
-Measurement creates values 0.76 9.61e-03 2.3
-Probability from uncertainty 0.39 3.24e-01 0.9
-Measurement disturbs system 0.97 1.91e-06 4.1
-Uncertainty linked to disturbance 0.71 9.61e-03 1.8
-Table 5: Degree of consensus from the graduate sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
-Assertion Affirmation $p$ $\sigma$
-Collapse 0.79 6.47e-03 2.1
-Non-ensemble psi 0.93 6.10e-05 3.2
-Ontologically complete 0.93 6.10e-05 3.2
-No trajectories 0.54 2.12e-01 0.3
-Complementarity 0.50 3.95e-01 0.0
-Wave-particle duality 0.71 2.87e-02 1.6
-Single particle uncertainty 0.82 9.16e-04 2.4
-Single particle wave-like 0.79 6.47e-03 2.1
-Measurement creates values 0.64 8.98e-02 1.1
-Probability from uncertainty 0.29 8.98e-02 1.6
-Measurement disturbs system 0.86 9.16e-04 2.7
-Uncertainty linked to disturbance 0.71 2.87e-02 1.6
-Table 6: Degree of consensus from the ‘Wikipedia notable’ sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
-Assertion Affirmation $p$ $\sigma$
-Collapse 0.74 2.45e-02 1.9
-Non-ensemble psi 0.94 7.63e-06 3.6
-Ontologically complete 0.88 1.37e-04 3.2
-No trajectories 0.35 1.66e-01 1.2
-Complementarity 0.41 3.15e-01 0.7
-Wave-particle duality 0.71 2.45e-02 1.7
-Single particle uncertainty 0.82 1.17e-03 2.7
-Single particle wave-like 0.76 6.36e-03 2.2
-Measurement creates values 0.74 2.45e-02 1.9
-Probability from uncertainty 0.32 1.66e-01 1.5
-Measurement disturbs system 0.94 7.63e-06 3.6
-Uncertainty linked to disturbance 0.79 1.17e-03 2.4
-Table 7: Degree of consensus from the ‘modern’ sample. $p$ is the probability of obtaining the given degree of consensus by chance from a random, binomial distribution.
-Assertion Affirmation $p$ $\sigma$
-Collapse 0.71 1.13e-02 2.0
-Non-ensemble psi 0.90 1.49e-06 3.9
-Ontologically complete 0.79 7.72e-04 2.9
-No trajectories 0.35 7.58e-02 1.4
-Complementarity 0.40 2.71e-01 1.0
-Wave-particle duality 0.71 1.13e-02 2.0
-Single particle uncertainty 0.60 1.54e-01 1.0
-Single particle wave-like 0.75 3.31e-03 2.4
-Measurement creates values 0.67 3.20e-02 1.6
-Probability from uncertainty 0.23 1.13e-02 2.7
-Measurement disturbs system 0.85 1.39e-04 3.5
-Uncertainty linked to disturbance 0.48 5.81e-01 0.2
-We begin with propositions related to the nature of the quantum state.
+We display these in terms of 4 samples: the full sample (black), undergraduate books (yellow), graduate texts (blue), works on the ‘Notable books’ list from Wikipedia (red), pre-1980 works (gray), post-1980 works (green), and post-2000 works (magenta).
+The only question substantially impacted by inference of response is “measurement” creates values.
+We supplement this plot with Fig.
+2, which shows that rejection is not inferred to any significant degree.
+It would be ideal to include the significance of the consensus in each case.
+However, there is no obvious statistical prescription for doing so, as responses to the propositions are not generated by random processes in any sense.
+We note however, that many positions here achieve agreement frequencies above 80%.
+Given a sample of $42$ works, this is sign of substantial consensus.
+Figure 1: Affirmation frequency for responses to each question.
+The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), pre-1980 (grey), post-1980 (green), and post-2000 (magenta).
+The error bars reflect the effect of including values inferred for logical consistency.
+Figure 2: Rejection frequency for responses to each question.
+The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), pre-1980 (grey), post-1980 (green), and post-2000 (magenta).
+The error bars reflect the effect of including values inferred for logical consistency.
+We begin by discussing propositions related to the nature of the quantum state.
 Interestingly, we see that not all the books regard $\psi$ as ontologically complete.
-However, the support for this proposition is robust with assent frequency $0.85$ ($p=5.46\times 10^{-6}$) in the full sample, with all of sub-samples having assent frequency $\geq 0.79$ ($p\leq 7.72\times 10^{-4}$).
-It is interesting that the lowest support for this occurs within the undergraduate and ‘modern’ samples.
+However, the support for this proposition is robust with assent frequency $0.85$ in the full sample, with all subsamples having assent frequency $\geq 0.8$.
+It is interesting that the lowest support for this occurs within the undergraduate and post-2000 samples, with outright rejection passing $0.1$ for post-1980 and post-2000 subsamples.
+Notably, there is no rejection in the pre-1980 sample, giving an indication that responses to this proposition are evolving.
 In terms of dissent, Phillips [32] and Moore [36] limit themselves to saying that $\psi$ can determine all the probabilities associated with a system but do not go on to say this is all the information allowed.
 Whereas, [21] openly suggests that interpretive difficulties might be overcome by finding a more general theory to supersede quantum mechanics.
-We also see strong support (a minimum frequency of $0.9$ with $p\leq 6.1\times 10^{-5}$) for the notion that a quantum state describes a single instance (as opposed to an ensemble).
-Notably, this achieves $0.95$ ($p=1.16\times 10^{-10}$) assent in the full sample and the undergraduate works are near unanimous on this count.
-Support is weaker for state collapse, where we obtain sub-sample assent frequencies above $0.71$ ($p\leq 2.45\times 10^{-2}$) while achieving $f=0.77$ ($p=1.62\times 10^{-4}$) in the full sample.
+We also see strong support (a minimum frequency of $0.94$) for the notion that a quantum state describes a single instance (as opposed to an ensemble).
+Notably, this achieves $0.96$ assent in the full sample and the undergraduate works are near unanimous on this count.
+In state collapse we find an interesting case.
+Here, both the assent and rejection frequencies increase as we move from pre-1980 to more modern subsamples.
+This seems to indicate the point is becoming more polarising over time.
+The collapse proposition still has subsample assent frequencies above $0.75$ while achieving $0.79$ in the full sample.
 From these results we can also see that an orthodox view requires the state to contain the entire ontological situation and apply to individual particles.
@@
 Of the three propositions, only affirmation of collapse reduces slightly if we consider sources after 1980.
-Thus, the orthodoxy on these issues seems largely unchanged over the last century.
-It is noteworthy that, in a sample of $33$ works, there are only two dissenters to the notion that quantum states collapse: Weinberg [17] as well as Binney & Skinner [18].
-Additionally, the topic is not mentioned by Griffiths, Peres, Müller-Kirsten, and Schiff [28, 29, 22, 46].
-Similarly, only Peres [29] dissents from the notion that quantum states describe individual instances.
+There are tentative signs of evolution in the consensus here.
+Particularly, the rate of rejection for collapse and ontological completeness both increase substantially when the post-1980 (and post-2000) subsample is compared to the full set.
+However, the assent frequency for collapse also increases from pre to post-1980, perhaps a marker of increased controversy.
+It is also noteworthy that, in a sample of $42$ works, there are only three dissenters to the notion that quantum states collapse: Weinberg [21], Merzbacher [51], as well as Binney & Skinner [22].
+Additionally, the topic is not mentioned by Peres, Mandl, Müller-Kirsten, and Schiff [38, 39, 29, 57].
+Only Peres [38] dissents from the notion that quantum states describe individual instances.
 Interestingly, we note that Home [6] argues that collapse is not essential to characterise the orthodoxy.
@@
 Now we turn to propositions on the nature of particles.
-A strong degree of consensus is clear around the notion that individual particles display wave-like behaviour with assent frequency $0.85$ at ($p\leq 5.46\times 10^{-6}$), sub-sample frequency $\geq 0.75$ ($p\leq 6.36\times 10^{-3}$).
+A strong degree of consensus is clear around the notion that individual particles display wave-like behaviour with assent frequency $0.88$ in the full sample, while maintaining subsample frequencies $\geq 0.78$.
 The only dissenters here being Weinberg [21], Binney & Skinner [22], and Sakurai [42].
 The topic is also not discussed in von Neumann [55] (perhaps unsurprising given the mathematical focus) and Griffiths [37].
-Reasonable support also exists for the notion of wave-particle duality with assent $0.79$ ($p\leq 1.62\times 10^{-4}$) in the full sample, sub-sample frequency $\geq 0.71$ ($p\leq 2.87\times 10^{-2}$).
-Here only von Neumann [44] and Binney & Skinner [18] outright denying this proposition.
-It is also not mentioned by Weinberg [17], Mandl [30], Townsend [31], and Sakurai [33].
+Interestingly, the lowered assent in the post-1980 subsample seems to stem from more recent graduate texts rejecting the proposition.
+There is also notable evolution, in the sense that there are no rejections pre-1980, with these increasing in frequency as we move to post-1980 and post-2000 subsamples.
+Reasonable support also exists for the notion of wave-particle duality with assent $0.78$ in the full sample and subsample frequency $\geq 0.71$.
+Here only von Neumann [55], Binney & Skinner [22], and Tong [16] outright deny this proposition.
+It is also not mentioned by Weinberg [21], Mandl [39], Townsend [40], Davydov [49] and Sakurai [42].
+There is discernable evolution based on publication year, rejection is much more common in the post-2000 subsample than either post-1980 or pre-1980 cases.
 For the proposition that individual particles being subject to the uncertainty principle, that is, one cannot know a given particle’s position and momentum (for example) simultaneously to arbitrary precision, we see more controversy.
-This is so because the full sample finds assent of $0.82$ ($p=3.31\times 10^{-5}$), however, the ‘modern’ subs-sample only registers $0.6$ ($p=0.154$).
-The reason for this change is that Griffiths [28], Peres [29], McIntyre [20], Moore [27], and Sakurai [33] all actively deny this proposition.
-This shows some evidence of a change in the consensus over time.
-However, a new orthodoxy is not yet entrenched, as Binney & Skinner [18], Longair [19], Townsend [31], Konishi & Paffuti [21], Omnés [26], Shankar [34], Phillips [24], Greiner [32], and Le Bellac [23] all affirm the proposition.
-Additionally, it is reasonably inferred that Mandl [30] and Weinberg [17] also support it.
-Home [4] does not include this in their conception of the orthodoxy on the grounds that is a flawed proposition.
+This is so because the full sample finds assent of $0.85$ and the pre-1980 subsample has nearly $1.0$.
+However, the post-1980 subsample only registers $0.77$, this then resurges to over $0.9$ in the post-2000 subset.
+This proposition receives the greatest rejection frequency in the post-1980 subsample.
+However, it is not rejected at all in the pre-1980 or post-2000 cases.
+The works between 1980 and 2000 that cause this are Griffiths [37], Peres [38], McIntyre [24], Moore [36], and Sakurai [42].
+This shows some evidence of a change in the consensus over time, but it appears to have been a fluctuation, returning towards the pre-1980 consensus after two decades.
+Home [6] does not include this proposition in their conception of the orthodoxy, likely as it is a flawed proposition.
 As argued in [9], the uncertainty principle could not function in the suggested manner.
 If the measurements from individual particles had the uncertainty principle’s scatter, the final ensemble distribution would have a larger uncertainty.
-Thus, preventing the verification of the principle as derived, where it applies to standard deviations over an ensemble of trials.
-Despite scepticism on this point, it is clearly wrong to exclude this from the orthodoxy, as the idea is uniformly affirmed by older works and still prevalent, but less so, in modern texts.
-It is notable that affirmation of complementarity is generally weak, even the full sample only finds an agreement frequency of $0.53$ ($p=0.243$).
-This is not due to denial but simply it never being mentioned in many later texts.
+Thus, preventing the verification of a lower bound.
+Despite scepticism on this point, it is clearly wrong to exclude this from the orthodoxy, as the idea is uniformly affirmed by older works and is resurgent in recent texts.
+It is possible that this could be connected to the rise of interpretations of QM that use “information” as an explanatory tool.
+A common formulation of the uncertainty principle in terms of “allowed information” then becomes very tempting.
+It is notable that affirmation of complementarity is generally weak, even the full sample only finds an agreement frequency of $0.52$.
+This is not due to denial (no one actually rejects it explicitly) but simply never being mentioned in many later texts.
+This accounts for the sustained drop in assent frequency from pre-1980 to post-1980 and post-2000 samples.
 Regardless, it cannot have a place in any robust orthodoxy.
-We also see that the issue of whether particles possess trajectories is contentious, with frequency $0.45$ ($p=0.364$) in the full sample.
-This becomes even more contentious in the ‘modern’ sample with $0.35$ ($p=0.0758$).
-This proposition still finds agreement in modern works like Konishi & Paffuti [21], Cohen-Tannoudji [36], Shankar [34], Greiner [32], and Phillips [24], with partial agreement in Moore [27] and Longair [19].
-However, the mode response is 00 in the undergraduate, ‘Wikipedia notable’ and ‘modern’ samples.
+We also see that the issue of whether particles possess trajectories has become contentious, with frequency $0.51$ in the full sample.
+This becomes even more contentious in the post-1980 and Wikipedia samples with $0.44$ and $0.39$ respectively.
+It initially enjoyed an assent frequency $\sim 0.6$ pre-1980 but has dropped to around $\sim 0.4$ in more recent works.
+The idea is never rejected pre-1980, but rejection appears afterwards and is largely maintained in more recent works.
+Asserting there are no trajectories still finds agreement in modern works like Kok [20], Miller [27], Jain [28], Basdevant [31], Konishi & Paffuti [26], Cohen-Tannoudji [45], Shankar [43], Greiner [41], and Phillips [32].
+There is also partial agreement in Moore [36] and Longair [23].
+However, the mode response is 0 in the undergraduate, ‘Wikipedia notable’, and post-1980 samples.
 So there is some indication that this proposition is increasingly marginalised by the orthodoxy.
@@
 Next, we turn to propositions around the question of why quantum mechanics predicts only statistical results.
-There is clear consensus, regardless of whether or not we confine our attention to modern works, that measurement disturbs quantum systems with assent $0.92$ ($p=6.54\times 10^{-8}$) in the full sample.
-Notably, the undergraduate sample has a frequency of $\sim 1.0$ for this proposition and only the graduate and ‘modern’ samples drop below $0.94$.
+There is clear consensus, regardless of whether we confine our attention to modern works, that measurement disturbs quantum systems with assent $0.91$ in the full sample, with a minimum of $0.81$ in the graduate case.
+Notably, the pre-1980 and post-2000 samples have very similar assent frequencies.
 Indeed, only Peres [38] actively dissents.
 Their dissent is justified by the possibility of non-interactive measurement in QM [59].
-There is also support for the idea that measurement itself actually creates the observed values with frequency $0.71$ ($p=2.28\times 10^{-3}$) in the full sample while retaining $0.67$ ($p\leq 3.2\times 10^{-2}$) in the ‘modern’ sub-sample.
-It should be noted that the affirmation is only stated openly 16 out of 29 times, despite being a plausible corollary to “measurement disturbs quantum systems” and that there are no active dissenters, not even [29].
-The low significance thus represents the number of times the assent was only inferred.
-Only three authors do not provide grounds for inferring agreement: Sakurai [33], Mandl [30], and Weinberg [17].
+There is also support for the idea that measurement itself actually creates the observed values with frequency $0.76$ (without inferred values this drops to $0.59$) in the full sample while retaining 0.730.73($0.58$) in the post-1980 subsample.
+Interestingly, the assent frequency drops further in the post-2000 case, indicating that there is some evolution on this proposition.
+However, it is never rejected in any sample.
+It should be noted that the affirmation is only stated openly around 6060% of the time, despite being a plausible corollary to “measurement disturbs quantum systems” and that there are no active dissenters, not even [38].Only three authors do not provide grounds for inferring agreement: Sakurai [42], Mandl [39], and Weinberg [21].
 The common case for inference is as follows: $\psi$ applies to individual systems, hence superpositions are real.
 Measurement disturbs the system and changes a superposition to an eigenstate, so it is either changing an unknown underlying value or creating a definite one where there was none.
-In either case the pre-existing value is totally superfluous if the collapse process is irreversible and indeterminist.
-The notion that probabilities in QM result from the uncertainty principle has always been a minority view it appears, as most authors do not mention anything like it (this question has a mode of 00 in all samples).
-There is still some support for the notion that the uncertainty principle is a consequence of measurement disturbance, with assent frequency $>0.71$ outside the ‘modern’ sample where it is only $0.48$.
-This indicates that affirmation of this proposition may be falling out of the orthodoxy.
-Interestingly, the sample with most affirmation for this last proposition is the “Wikipedia notable” one, with $0.79$ assent ($p=1.17\times 10^{-3}$).
+In either case the pre-existing value is totally superfluous if the collapse process is irreversible and indeterminate.
+However, the statistics here are significantly affected by the reliability of the inferences being made.
+The notion that probabilities in QM result from the uncertainty principle has always been a minority view it appears, as most authors do not mention anything like it, achieving a maximum assent of $0.4$ in the pre-1980 sample.
+However, it is not rejected until after 1980 with the rejection rate dropping substantially in the post-2000 works.
+This coincides with an increase in assent post-2000, where Tong [16], Tamvakis [19], Longair [23], Konishi & Paffuti [26], and Müller-Kirsten [29] all support the position.
+In the same sample, only Binney & Skinner [22] reject it, so the mode response is still to leave it out.
+There is still some support for the notion that the uncertainty principle is a consequence of measurement disturbance, with assent frequency $0.69$.
+It falls from $0.8$ pre-1980 to around $0.6$ post-1980 and 2000.
+Thus, providing clear evidence for consensus evolution on this point.
+Interestingly, the sample with most affirmation for this last proposition is the “Wikipedia notable” one, with $0.79$ assent.
+The only outright rejections here are Rae [15] and Basdevant [31], who both note the uncertainties can happily be measured in separate runs of an experiment, so “disturbance” need not occur.
 Thus, for the orthodoxy, we do not reveal the values of studied quantum systems but force them to assume definite values (or alter them) via some uncontrolled disturbance attached to the abstract notion of measurement.
 The “uncontrolled” aspect of this is vital to square the statistical nature of QM with the proposition that quantum states apply to individual instances of systems.
-Here we see emergence of anti-realism as a pillar of the orthodoxy.
-This is entailed by the fact that we are not exploring the pre-existing properties of quantum systems so much as forcing them into altered states under measurement.
-To better understand the development of the orthodoxy over time we can also consult the mean responses in each sub-sample, as in Figure 2.
-Here we can see the modern sample has a consistently lower mean response to the vast majority of propositions.
-The most significant difference occurs for the “single particle uncertainty” proposition, with a $40$% reduction in the mean value relative to the other samples.
-This reduction is twice the difference in assent frequency between the samples, indicating the presence of active dissent from the proposition.
-This cannot be seen from the assent frequency, as $-1$ and 00 have the same effect on the frequency of $1$.
-The proposition of “ontological completeness” sees a decrease in mean response of around 20% in the ‘modern’ sub-sample that is not evident in response frequency, again showing evidence of reduced consensus on this point.
-The “uncertainty from disturbance” proposition also has a mean response decline of $30$%.
-However, this correlates well with the assent frequency, indicating only that this idea is increasingly not mentioned in more recent works.
-These demonstrate that there are propositions that were previously orthodox (with assent frequencies above $0.7$ in other samples) that have become more controversial over time.
-Some less popular positions (“probability from uncertainty” and “no trajectories”) also exhibit evidence of increased controversy in the modern sample, showing smaller mean values than even other samples where the mode is similarly 00.
-Figure 2: Average values of the responses to each question. The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), and modern (grey).
+Next, we can ask whether there are correlations between proposition responses.
+To do this, we determine the frequency with which answers for different pairs of propositions match.
+We present the results in Table 3 for the full and post-1980 samples of textbooks.
+We choose only to display those for which the assent frequency exceeds $0.6$ to determine if there is a strong correlation between orthodox positions.
+Importantly, the strongest correlation occurs for the two pillars of the orthodoxy: “measurement disturbs systems” and “non-ensemble $\psi$”, indicating that they are correlated 9595% of the time.
+Over all, it is evident that relatively strong pair-wise correlations exist between all the propositions with assent frequency above $0.6$.
+They all correlate strongly to one (or both) of the two pillar propositions as well.
+This demonstrates that selecting an orthodoxy on this basis yields a consistently affirmed set of propositions.
+Proposition 1 Proposition 2 Full sample Post-1980
+Collapse Non-ensemble $\psi$ 0.81(0.79) 0.81(0.81)
+Collapse Ontologically complete 0.76(0.74) 0.74(0.74)
+Collapse Wave-particle duality 0.69(0.67) 0.74(0.74)
+Collapse Single particle uncertainty 0.71(0.69) 0.67(0.67)
+Collapse Single particle wave-like 0.79(0.76) 0.81(0.81)
+Collapse Measurement creates values 0.81(0.52) 0.81(0.56)
+Collapse Measurement disturbs system 0.81(0.76) 0.81(0.78)
+Non-ensemble $\psi$ Ontologically complete 0.83(0.81) 0.81(0.78)
+Non-ensemble $\psi$ Wave-particle duality 0.74(0.74) 0.70(0.70)
+Non-ensemble $\psi$ Single particle uncertainty 0.90(0.88) 0.85(0.81)
+Non-ensemble $\psi$ Single particle wave-like 0.86(0.83) 0.81(0.78)
+Non-ensemble $\psi$ Measurement creates values 0.90(0.57) 0.85(0.56)
+Non-ensemble $\psi$ Measurement disturbs system 0.95(0.90) 0.96(0.89)
+Ontologically complete Wave-particle duality 0.64(0.64) 0.63(0.63)
+Ontologically complete Single particle uncertainty 0.79(0.76) 0.74(0.70)
+Ontologically complete Single particle wave-like 0.79(0.79) 0.78(0.78)
+Ontologically complete Measurement creates values 0.81(0.52) 0.78(0.48)
+Ontologically complete Measurement disturbs system 0.81(0.79) 0.81(0.78)
+Wave-particle duality Single particle uncertainty 0.67(0.67) 0.59(0.59)
+Wave-particle duality Single particle wave-like 0.76(0.76) 0.74(0.74)
+Wave-particle duality Measurement creates values 0.83(0.55) 0.85(0.63)
+Wave-particle duality Measurement disturbs system 0.74(0.71) 0.74(0.70)
+Single particle uncertainty Single particle wave-like 0.83(0.81) 0.78(0.74)
+Single particle uncertainty Measurement creates values 0.83(0.50) 0.74(0.44)
+Single particle uncertainty Measurement disturbs system 0.86(0.83) 0.81(0.78)
+Single particle wave-like Measurement creates values 0.86(0.55) 0.81(0.56)
+Single particle wave-like Measurement disturbs system 0.83(0.81) 0.81(0.78)
+Measurement creates values Measurement disturbs system 0.90(0.55) 0.89(0.56)
+Table 3: The frequency with which answers for two propositions match in the full and post-1980 samples. The values in brackets reflect the case without inference of responses.
 Finally, we have Figure 3. This displays the “orthodoxity” of each book, i.e. how close they are to the averaged responses over the full sample. This is determined following
@@
 where the sum runs over the propositions, $w_{i}$ is the average response value, and $a_{i}$ is that of the reference work in question.
-Note that we exclude propositions that do not have assent frequency $>0.6$.
-The average orthodoxity is $74\pm 25$%, indicating a robust degree of agreement on average.
-For convenience we order Fig.
+Note that we exclude propositions that do not have a full-sample assent frequency >0.6>0.6.
+As illustrated above, this selection criterion results in set of well-correlated propositions.
+The average orthodoxity is 7474% with standard deviation 2323% indicating a robust degree of agreement on average.
+This includes inferred values, but, this effect is small, going to 7171% and 2424% without them.
+Note that these averages are robust to increasing the frequency inclusion threshold to $0.7$ as well.
+For convenience, we order Fig.
 3 by year of first publication.
 This allows us to determine if there are any trends in the orthodoxy.
-What we see is quite remarkable.
-Early works adhere very closely to the orthodoxy (as might be expected).
+Early works adhere very closely to the full-sample orthodoxy, with those before 1980 having an average orthodoxity of 8383% with standard deviation 1414%.
 The period around the early 90’s shows a fluctuation of far lower agreement with the orthodoxy.
-However, the orthodoxy reasserts itself after this point (but slightly more weakly than in the earlier works).
+This results in post 1980 works having an average orthodoxity of 6868% and standard deviation 2525%.
+The substantial nature of this change from the earlier consensus is most vividly illustrated by the 50% increase in the post-1980 standard deviation, despite a 7373% increase in sample size.
 The 90’s dip in the orthodoxity is constituted by the works: Peres [38], Moore [36], Griffiths [37], Mandl [39], Rae [15], and Sakurai [42].
-Overall, there is some evidence for an evolutionary trend in the orthodoxy, but it must be admitted that modern works quite closely follow the interpretation put forward by their predecessors.
-Despite this as an overall statement, we can see recent dissent from Weinberg [17] and some clearly unorthodox positions in both Binney & Skinner [18] and Griffiths [28].
-This is of interest given the prominence of the first’s author and that the latter two also appear on the “Wikipedia Notable” list.
-It is also interesting that the modern reduction in orthodoxity (even if unspectacular on average) revolves around the questions of the wave-like behaviour of individual particles.
-Something that is denied by Weinberg, Binney & Skinner, Griffiths, and Sakurai [17, 18, 28, 33], these being the works with some of the lowest orthodoxity.
-The lack of orthodoxity for Peres [29] is rooted in the author’s instrumentalism (being far closer to the version of a Copenhagen viewpoint as understood by Stapp [2]).
-For additional works that were not included in the sample we find the following orthodoxity measures: Ballentine [12] $-0.89$, Home [4] 0.44, Hannabuss [14] 0.43, and Nielsen & Chuang [15] 0.44.
+Of these, Mandl is here mainly due to omission, rather than contesting orthodox positions.
+If we replace Mandl with Weinberg [21], we form a set of all the authors who deny more than one proposition.
+However, there is no consistent pattern in these denials.
+Overall, there appears to be a post-1980 divergence from the earlier orthodoxy.
+This being said, it resembles a fragmentation, rather than a new consensus emerging, due to the lack of consistent patterns in the denial of previously orthodox positions.
+However, it must be admitted that many modern works adhere quite closely to the interpretation put forward by their predecessors.
+This suggests the possibility that our present epoch is the start of a collapse in confidence in the orthodoxy.
+A survey of working physicists has recently demonstrated that the community is fragmented on the interpretation of quantum mechanics [2], notably more so than a previous survey in 2016 [1].
+This could well signal an increase in divergence from the previous quantum orthodoxy in future textbooks.
+Somewhat complicating such an idea is the tendency of more orthodox post-1980 works to cite Bell’s theorem as a vindication of many aspects of the orthodoxy, perhaps accounting for the apparent resurgence in orthodoxity after the late 1990’s.
+The timeline notably aligns with the emergence of “Copenhagen+” interpretations like quantum Bayesianism (see [60]).
+For additional works that were not included in the sample we find the following orthodoxity measures: Ballentine [14] 0.88-0.88, Home [6] 0.42, Hannabuss [17] 0.45, and Nielsen & Chuang [18] 0.42.
 The last of these, Nielsen & Chuang [18], only shows a lower orthodoxity by virtue of never mentioning the topic of waves.
 Their strong endorsement of Peres [38] and Sakurai [42] might indicate little sympathy with this aspect of the orthodoxy, but the clarity of this is marred by also endorsing Cohen-Tannoudji [45].
-Interestingly, Home’s characterisation of the orthodoxy [4] appears incomplete at $0.44$.
-Additionally, this is only when one concentrates on the aspects attributed to the non-ensemble case.
+Interestingly, Home’s characterisation of the orthodoxy [6] appears incomplete at $0.42$.
+Additionally, this is only when one concentrates on the aspects they attributed to the non-ensemble case.
 Home considers ensemble views as part of the orthodoxy on the basis that they view the formalism as being complete (there are no supplemental entities needed to explain what is going on).
@@
 Since Copenhagen-like interpretations hinge so heavily on collapse, their popularity implies that Home’s primary condition of orthodoxy becomes more doubtful.
-Interestingly, Ballentine [12], who would express an orthodox view according to Home, achieves an orthodoxity of $-0.89$ here, dissenting from (or not addressing) every proposition considered.
+Interestingly, Ballentine [14], who would express an orthodox view according to Home, achieves an orthodoxity of 0.88-0.88 here, dissenting from (or not addressing) every proposition considered.
 This is because Home’s rationale for including ensemble views in the orthodoxy may not be well justified.
@@
 Figure 3: The “orthodoxity” or degree to which each studied source agrees with the averaged responses to each proposition. This is ordered by date of first publication.
+We can study how time period effects the orthodoxy by change the orthodoxy weights to those of a subsample.
+We display the changes in orthodoxy for both pre and post-1980 selections in Fig.
+4.
+Importantly, the pre-1980 case includes two extra propositions: “no trajectories” and “complementarity”, as both pass the $0.6$ assent threshold in that subsample.
+For both our modified orthodoxies, the mean changes from the full sample are small, being $0.0049$ and $0.015$ respectively.
+This indicates a small preference for the pre-1980 orthodoxy, with several authors becoming around 1010% more orthodox in this case.
+In contrast, the post-1980 case has smaller orthodoxity gains and several large decrements.
+Interestingly, the largest post-1980 decreases are all for post-1980 works.
+Whereas, the largest gains for a pre-1980 orthodoxy are from post-1980 works as well.
+This indicates that the most robust orthodoxy can be drawn from pre-1980 works’ positions, but there are definite hints of evolution.
+However, there are no significant correlations between the changes (for any of our temporal subsamples), indicating that a new consistent orthodoxy is yet to emerge.
+Figure 4: The changes in “orthodoxity” that result if we use different samples to define the orthodoxy.
 We can now ask how far the revealed orthodoxy diverges from the idea that it can be defined entirely by the ontological primacy of the quantum state (as advocated by Home [6]).
@@
 This is a vital part of the orthodoxy (stronger even than ontological completeness) and it does not follow from $\psi$ being ontologically complete.
-There being no reason to preclude a fundamentally indeterminist viewpoint which does not require single realisations to be associated to $\psi$.
-If this is the case then there are a cascade of propositions that must follow, as they all depend on “non-ensemble $\psi$” for their relevance.
+There being no reason to preclude a fundamentally indeterminist viewpoint.
+If this is the case, then there are a cascade of propositions that must follow, as they all depend on “non-ensemble $\psi$” for their relevance.
 These are “single particles are wave-like”, “states collapse on measurement”, “measurement inherently disturbs the system”, and “measurement creates observed values”.
@@
 6 Conclusion
-When we consider a sample of $33$ textbooks, aimed at both graduate and undergraduate students of quantum mechanics, we find a robust orthodoxy emerging.
-This is powerfully illustrated by the average orthodoxity of the sources being $74$% with a standard deviation of only $25$%.
-Given that the studied sample spans nearly a century, this provides good evidence for a mostly unchanged consensus on the interpretation of quantum mechanics.
-In modern works there is slightly less support for the notion that individual particles are wave-like, but large reductions in support for the notions that the uncertainty principle applies to individual particles, and that quantum uncertainty results from measurement disturbance.
+When we consider a sample of $42$ textbooks, aimed at both graduate and undergraduate students of quantum mechanics, we find a robust orthodoxy emerging.
+This is powerfully illustrated by the average orthodoxity of the sources being 7373% with a standard deviation of 2424%.
+Given that the studied sample spans nearly a century, this provides good evidence that many aspects of the consensus have remained unchanged.
+In post-1980 works, the average orthodoxity is reduced to 6868% from a pre-1980 value of 8484%.
+In particular, there is less support for the notions that individual particles are wave-like, and that the quantum state is ontologically complete.
+There are also more significant reductions in support for the notions that the uncertainty principle applies to individual particles, quantum uncertainty results from measurement disturbance, particles do not have trajectories, and complementarity.
+The reduction in orthodoxy over time mirrors recent surveys of physicists attitudes in quantum mechanics [2].
+Perhaps indicating the start of opinion fragmenting away from the orthodoxy, as there are no obvious patterns for groups of propositions being denied together.
 The revealed orthodoxy strongly resembles many aspects of Copenhagen-like frameworks.
-Particularly, there is powerful agreement (assent frequency $0.92$) that measurement inherently disturbs quantum systems, single instances of systems are individually described by a quantum state, rather than as an ensemble (assent frequency $0.95$), state collapse happens (assent frequency $0.77$) , that individual particles are wave-like (assent frequency $0.85$), and that quantum states are ontologically complete (assent frequency $0.85$).
-These positions are further supplemented by the existence of a wave-particle duality (assent frequency $0.79$) and that measurement actually creates or alters the definite property values being measured (assent frequency $0.71$).
-More contentious, but still popular, are the notions that the uncertainty principle arises from measurement disturbance (assent frequency $0.71$), and that individual particle measurements are subject to the uncertainty principle (assent frequency $0.82$).
-These assent frequencies have a very small probability to arise by chance when drawing samples from a random binomial distribution, indicating a level of statistical robustness in the observed consensus.
+Particularly, there is powerful agreement (assent frequency $0.91$) that measurement inherently disturbs quantum systems, single instances of systems are individually described by a quantum state, rather than as an ensemble (assent frequency $0.96$), state collapse happens (assent frequency $0.79$), that individual particles are wave-like (assent frequency $0.88$), and that quantum states are ontologically complete (assent frequency $0.85$).
+These positions are further supplemented by the existence of a wave-particle duality (assent frequency $0.78$) and that measurement actually creates or alters the definite property values being measured (assent frequency $0.76$, but only $0.59$ without inferred values).
+More contentious, but still popular, are the notions that the uncertainty principle arises from measurement disturbance (assent frequency $0.67$), and that individual particle measurements are subject to the uncertainty principle (assent frequency $0.85$).
 It is clear that the orthodoxy favours the view of the quantum state as a complete description of the actual state of affairs, rather than the more instrumental view attributed to Copenhagen by authors like Stapp [4].
 In this regard, the views of the orthodoxy espoused by Home [6] and Jammer [5] are vindicated.
-The work most embodying this kind of orthodoxy would be Peres [29], with similarities in Mandl [30], Nielsen & Chuang [15], Hannabuss [14], and Sakurai [33].
-However, it is not clear that all the logical consequences of the single proposition view are believed by the physics community (and it is clear there are wave-related addenda).
+The work most embodying Stapp’s version would be Peres [38], with similarities in Mandl [39], Nielsen & Chuang [18], Hannabuss [17], and Sakurai [42].
+However, it is not clear that all the logical consequences of Home’s view (that the orthodoxy is entirely characterised by ontological completeness) are believed by the physics community (and it is clear there are wave-related addenda).
 Thus, the orthodoxy’s characterisation requires more nuance than being distilled to a single point about the status of quantum states.
-It is clear that, examining all the propositions in this work, there is no room in the orthodoxy for ensemble views, entirely in opposition to the position of Home.
+It is clear that, examining all the propositions in this work, there is no room in the orthodoxy for ensemble views, entirely in opposition to the stance taken by Home.
 In fact, it is a strong requirement of the orthodoxy that quantum states describe single instances, rather than ensembles.
+It is also interesting to note that respondents in [1] characterise the Copenhagen interpretation’s most salient features as “collapse” (77%), “complementarity” (71%), “indeterminism” (46%), and “correspondence” (43%).
+This is quite different to the orthodoxy deduced from our sample of textbooks.
+Where complementarity is far less important, indeterminism is more strongly supported (non-ensemble $\psi$ + ontological completeness), and collapse is popular but not absolutely primary.
+This suggests that, although physicists might predominantly report subscribing to the Copenhagen interpretation, the prevailing community beliefs do not fully conform to their own characterisation of this viewpoint.
 To an orthodox quantum mechanic, individual particles exist in wave-like superpositions, when left to their own devices, and assume their observed properties (and appear as particles) when forced to by outside interference.
 When this happens, the individual particle’s properties are subject to the uncertainty principle as a form of measurement uncertainty and the state is inevitably changed by the uncontrolled disturbance inherent in measurement.
-Perhaps understandably, few textbook authors state the anti-realist aspects of these views with complete openness.
-For instance the “measurements create observed values” proposition is explicitly endorsed by only around 50% of the texts but remains a plausible corollary of positions taken by another 40%.
+[…]
+Acknowledgements GB acknowledges funding from the National Research Foundation of South Africa under the incentive funding for rated researchers program.
+[…]
+Declarations Data used in this work is available on reasonable request. The authors declare no conflicts of interest.

### v3

@@
 Rather than surveying physicists, or poring over the precise details of the thoughts of Bohr and Heisenberg, we review a collection of 42 textbooks on quantum mechanics, encompassing the most popular and prominent works of this nature.
-We then gauge their response to 12 propositions to build up a picture of exactly what is believed by an orthodox quantum mechanic.
+We then gauge their response to 13 propositions to build up a picture of exactly what is believed by an orthodox quantum mechanic.
 We demonstrate that this orthodoxy has many aspects of Copenhagen-like viewpoints, but also shows some interesting emerging deviations.
 Moreover, it is more nuanced than some reductive characterisations that condense the orthodoxy down to the ontological primacy of the quantum state.
-The revealed orthodoxy has two consistent pillars: measurement inherently disturbs quantum states and these states refer to individual instances, not ensembles.
+The revealed orthodoxy has three consistent pillars: measurement inherently disturbs quantum states, these states refer to individual instances, not ensembles, and quantum systems do not have definite properties prior to measurement.
 More fully, it entails that individual particles exist in wave-like super-positions and present particle behaviours only when forced to by outside influences.
@@
 How do scientists understand quantum mechanics?
-The most popular interpretation of quantum mechanics (QM) is usually found in surveys to be “Copenhagen interpretation” [1, 2].
-However, the definition of this viewpoint is notoriously vague [3, 4, 5, 6], its characterisation tending to descend into hyperfine readings of Bohr and Heisenberg [4].
-Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [5].
+The most popular interpretation of quantum mechanics (QM) is usually found in surveys to be “Copenhagen interpretation” [sivasundaram2016surveyingattitudesphysicistsconcerning, survey2025].
+However, the definition of this viewpoint is notoriously vague [hanson1959, stapp1972, Jammer1974-JAMTPO-10, Home1997], its characterisation tending to descend into hyperfine readings of Bohr and Heisenberg [stapp1972].
+Max Jammer sets out the quantum formalism as a set of 5 postulates, all of which are necessary to an orthodox or Copenhagen-like view [Jammer1974-JAMTPO-10].
 These include the completeness of the state and the projection postulate.
-To produce something like a Copenhagen interpretation, this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [7] described as “doing epistemology without ontology”.
+To produce something like a Copenhagen interpretation, this is supplemented by a kind of positivist objection to ascribing as real anything that cannot be directly measured, generating what Shimony [shimony1993] described as “doing epistemology without ontology”.
 This all results in a similar view to Home, that it is the primacy of the quantum state that typifies the orthodox view.
-Interestingly, Stapp argues [4] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities.
+Interestingly, Stapp argues [stapp1972] that a Copenhagen viewpoint does not see the quantum state as describing the evolution of the actual system in question, only its probabilities.
 This seems somewhat difficult to square with the notion that the state is the most complete specification of the real situation, without some explanation of why additional information is inaccessible.
-Home [6] argues that the primary characteristic of an orthodox viewpoint is that the state is framed as a complete description of a quantum system.
+Home [Home1997] argues that the primary characteristic of an orthodox viewpoint is that the state is framed as a complete description of a quantum system.
 Thus, this encapsulates both Copenhagen-like and ensembles-based theories.
 This latter inclusion stems from a claim that ensemble theories are internally incomplete unless they assume completeness of the quantum state (otherwise they must supply some hidden variables to assign definite values to measurable properties).
-Home refers to Feyeraband’s encapsulation of the standard position as being that quantum systems possess no definite properties over and above those that can be derived from the state itself [8].
-These points reveal a tension in descriptions of a quantum orthodoxy: a Copenhagen-like view wants to claim a “completeness” or “finality” for QM (the “end of the road hypothesis” as Popper terms it [9]).
+Home refers to Feyeraband’s encapsulation of the standard position as being that quantum systems possess no definite properties over and above those that can be derived from the state itself [feyeraband1962].
+These points reveal a tension in descriptions of a quantum orthodoxy: a Copenhagen-like view wants to claim a “completeness” or “finality” for QM (the “end of the road hypothesis” as Popper termed it [Popper1982-POPQTA]).
 However, it also contains a strong degree of instrumentalism, as acknowledged by both Home and Jammer.
@@
 Rather than characterise this Copenhagen model, it seems more productive to instead determine what the consensus within the physics community actually is.
-With a particular focus on answers to questions about quantum phenomena, along the lines of those used in [1].
-Existing surveys of physicists tend to reveal a very fragmented community [1, 2], instead we will look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
+With a particular focus on answers to questions about quantum phenomena, along the lines of those used in [sivasundaram2016surveyingattitudesphysicistsconcerning].
+Existing surveys of physicists tend to reveal a very fragmented community [sivasundaram2016surveyingattitudesphysicistsconcerning, survey2025], instead we will look for a quantum mechanical orthodoxy in the many treatises that have followed from the foundational works usually analysed to define a Copenhagen interpretation.
 The target of our survey will be textbooks.
@@
 However, one suspects that such media would itself be based on textbooks, rather than wade through the mire of competing interpretations in the active literature.
-An aspect of this can be seen in the fact that more recent textbooks devote more space to the interpretation of quantum phenomena (many have this as a separate chapter even).
+An aspect of this can be seen in the fact that more recent textbooks devote more space to the interpretation of quantum phenomena (some have this as a separate chapter even).
 In addition, while working physicists may later deviate from the interpretations presented in reference works, we must acknowledge that their trajectory will be influenced by the initial conditions.
@@
 This is inherited by the vast majority of modern unorthodox interpretations (e.g.
-Bohmian mechanics [10], quantum Bayesianism [11], consistent histories [12], and relational quantum mechanics [13]), despite it not being necessary to the formalism [14].
+Bohmian mechanics [bohm1982], quantum Bayesianism [Fuchs_2013], consistent histories [sep-qm-consistent-histories], and relational quantum mechanics [Rovelli_1996]), despite it not being necessary to the formalism [ballentine1998quantum].
 Thus, even the beliefs of established researchers in quantum interpretation may be strongly correlated with the orthodoxy found in textbooks.
@@
 Once again, this injects subjectivity into our analysis, as it is not clear how we might choose these positions.
-We settle on a set of 12 propositions, divided into three categories.
-These are selected to at least cover the major assertions of a Copenhagen-like framework and show strong overlap with surveys [1].
+We settle on a set of 13 propositions, divided into three categories.
+These are selected to at least cover the major assertions of a Copenhagen-like framework and show strong overlap with surveys [sivasundaram2016surveyingattitudesphysicistsconcerning].
 The first category explores how we understand the quantum state, with the propositions: “the state is ontologically complete”, “the state does not merely describe an ensemble of identically prepared systems”, and “the state collapses upon measurement”.
-The second category revolves around the nature of particles with the propositions: “individual particles have wave-like properties”, “particle-like entities are described by a wave-particle duality”, “particles cannot simultaneously possess definite values of non-commuting observables”, “particles do not follow classical-like trajectories”, and “there is no adequate single picture of particle or wave”.
-Finally, we turn to the question of why quantum mechanics predicts statistics only.
-Here we have the propositions: “measurement inherently disturbs quantum systems”, “measurement creates the observed definite values”, “the uncertainty principle results from measurement disturbance”, and “the limitation to prediction of probabilities results from the uncertainty principle”.
+The second category revolves around the nature of particles with the propositions: “individual particles have wave-like properties”, “fundamental wave-like behaviour is reflected by the wave-particle duality”, “particles cannot simultaneously possess definite values of non-commuting observables”, “particles do not follow classical-like trajectories”, and “there is no adequate single picture of particle or wave”.
+Next, we turn to the question of why quantum mechanics predicts statistics only.
+Here we have the propositions: “measurement inherently disturbs quantum systems”, “quantum systems have no property values till measured”, “the uncertainty principle results from measurement disturbance”, and “the limitation to prediction of probabilities results from the uncertainty principle”.
+Finally, we have “other interpretations of quantum mechanics are irrelevant”, which doesn’t fall into any of the categories.
 These will all be elaborated on in later sections of this paper.
@@
 Defining all aspects of an interpretative orthodoxy is a difficult proposition.
-The approach taken here is to define three sub-categories of hopefully fundamental aspects of the interpretation of the quantum formalism, these being the nature of the quantum state, that of particles, and why our predictions are all statistical.
+The approach taken here is to define three sub-categories of hopefully fundamental aspects of the interpretation of the quantum formalism, these being the nature of the quantum state, that of particles, and why our predictions are probabilistic.
 Within each category we include propositions that source works can either agree with or dissent from.
@@
 It is entirely possible that this question set is incomplete (or contains some redundancy).
-We also note that our choices cover questions asked of working physicists in surveys [1, 2].
-Notably, the aforementioned [1] also includes “are values are defined prior to measurement?” (“measurement creates values”), “is randomness inherent in QM?” (strong overlap with “ontological completeness”), “where is the electron in a hydrogen orbital?” (“no trajectories”), “is macroscopic superposition possible?” (linked to “the state doesn’t just describe an ensemble”), and a question on the solution of the measurement problem is covered by the entire set of questions on why QM is statistical.
+We also note that our choices cover questions asked of working physicists in surveys [sivasundaram2016surveyingattitudesphysicistsconcerning, survey2025].
+Notably, the aforementioned survey [sivasundaram2016surveyingattitudesphysicistsconcerning] also includes “are values are defined prior to measurement?” (“no pre-measurement values”), “is randomness inherent in QM?” (strong overlap with “ontological completeness”), “where is the electron in a hydrogen orbital?” (“no trajectories”), “is macroscopic superposition possible?” (linked to “the state doesn’t just describe an ensemble”), and a question on the solution of the measurement problem is covered by the entire set of questions on why QM is probabilistic.
 However, we do not include questions about the role of the observer or the implications of Bell inequality violations.
@@
 The cases of incomplete overlap are down to the luxury of expositive works like textbooks.
-In that, our questions can be simpler, more numerous, and less nuanced than those in [1, 2], as they need not be processed in real time by the respondent.
+In that, our questions can be simpler, more numerous, and less nuanced than those in [sivasundaram2016surveyingattitudesphysicistsconcerning, survey2025], as they need not be processed in real time by the respondent.
 2.1 What is the quantum state?
@@
 The waves do not merely appear in the statistics of quantum systems.
-Viewpoints that might deny this proposition would be a de Broglie-Bohm interpretation [10] (where wave and particle co-exist as linked entities) or an ensemble one (where the waves are merely statistical patterns).
+Viewpoints that might deny this proposition would be a de Broglie-Bohm interpretation [bohm1982] (where wave and particle co-exist as linked entities) or an ensemble one (where the waves are merely statistical patterns).
 2.2.2 Wave-particle duality is necessary
@@
 More generally: systems cannot possess definite values of non-commuting observables simultaneously.
-This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble of measurements performed on similarly prepared systems [9].
-2.3 Why are our predictions statistical?
+This is opposed to a statistical understanding which recognises that the uncertainty principle is derived via standard deviations over an ensemble of measurements performed on similarly prepared systems [Popper1982-POPQTA].
+2.3 Why are our predictions probabilistic?
 When we conduct quantum mechanical experiments we require measuring over an ensemble of similarly prepared systems in order to compare to the probabilistic predictions of the theory.
@@
 Measurement inherently changes the state of a system while measuring it (except in the case of eigenstates).
-A viewpoint that denies this would note that non-interactive measurements exist in QM or that measurement has no abstract definition.
+A viewpoint that denies this would note that non-interactive measurements exist in QM, or that measurement has no abstract definition.
 2.3.2 Uncertainty results from disturbance
@@
 Particularly, this also requires single-particle uncertainty.
-The consequences of this is that measurements have probabilistic results, even though quantum mechanics is not formulated as fundamentally statistical.
+The consequences of this is that measurements have probabilistic results, even though quantum mechanics is not formulated as fundamentally probabilistic.
 One might deny this if one recognises that the uncertainty principle is a consequence of quantum rules, rather than a postulate, or if one notes the uncertainty principle is formulated in terms of statistical observables only.
-2.3.4 Measurement creates observed values
-The observed values of quantum properties are the result of an interaction between apparatus and system that alters the measured property.
-In this sense, they do not reflect underlying pre-measurement values in cases that are not eigenstates of the observable being measured.
-This stands in opposition to the usual scientific idea of measurement: that we are generally discovering the values of pre-existing properties.
-This assertion can take two forms: a strong form, where it is asserted that quantum systems do not actually possess definite properties till measured, and a weaker form where measurement merely changes pre-existing values.
-Both of these result in the pre-measurement value being indeterminate and potentially irrelevant.
+2.3.4 There are no pre-measurement values
+Except for those observable eigenstates, quantum systems do not have definite properties that we uncover via measurement.
+Instead, the value comes to exist at the point of measurement.
+A primary signature of this is how the Born rule is described.
+If a text states that $|\psi(x)|^{2}$ is the probability of finding the system with property value $x$, then it affirms this proposition.
+However, this alone is not enough, we only infer assent if there is other supportive text as well.
+A dissenting text would instead state the Born rule as giving the probability of a system having property value $x$.
+This appears to be a rather fine point of semantics.
+However, almost all the surveyed texts use precisely the same wording: finding, not having.
+The ones that don’t use “finding”, seem to do so deliberately, and it correlates with more direct assertions that pre-measurement values exist.
+2.4 Other propositions
+These don’t fit into the above categories.
+2.4.1 There is only the orthodoxy
+Only the orthodox presentation of quantum mechanics is worthwhile or serious.
+An affirming text either mentions no other interpretations or dismisses them all as unconvincing or disproven.
+A textbook that notes it is presenting the orthodox/conventional/standard approach but admits this isn’t definitive will be taken as dissenting.
 3 Sample of studied works
@@
 Note that Omnés work is included, even though it is a textbook that advocates against the orthodoxy in favour of “consistent histories”, as it provides its own description of the orthodox positions which we use for our study.
-In the case of [15], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the “conventional” approach.
-Note that page references for [16] are in reference to the online lecture notes on which it is based [^3] https://www.damtp.cam.ac.uk/user/tong/books/quantum.html, due to the recency of the text.
-We will also make reference to a set of 4 works whom we do not include in any of the samples: [6, 17, 18, 14].
+In the case of [rae2002], this thoughtfully discusses some foundational issues without expressing a direct opinion, here we substitute what the author characterises as the “conventional” approach.
+We will also make reference to a set of 4 works whom we do not include in any of the samples: [Home1997, Hannabuss:1999fa, nielsen2010, ballentine1998quantum].
 The first is excluded as it is a discussion of interpretations, but will provide a useful point of comparison.
@@
 Title Author Year Ref Level Wiki
-Lecture notes on quantum mechanics D. Tong 2025 [16] U N
-Basic quantum mechanics K. Tamvakis 2019 [19] U N
-A first introduction to quantum physics P. Kok 2018 [20] U N
-Lectures on quantum mechanics S. Weinberg 2013 [21] G N
-The physics of quantum mechanics J. Binney & D. Skinner 2013 [22] U Y
-Quantum concepts in physics M. Longair 2013 [23] G N
-Quantum mechanics D. McIntyre et al 2012 [24] U Y
-Quantum mechanics: theory and experiment M. Beck 2012 [25] U N
-Quantum mechanics: a new introduction K. Konishi & G. Paffuti 2009 [26] U N
-Quantum mechanics for scientists and engineers D. Miller 2008 [27] U N
-Quantum mechanics: a textbook for undergraduates M. Jain 2007 [28] U N
-Introduction to quantum mechanics H. Müller-Kirsten 2006 [29] U Y
-Quantum physics M. Le Bellac 2006 [30] G N
-Lectures on quantum mechanics J. Basdevant 2005 [31] G N
-Introduction to quantum mechanics A. Philips 2003 [32] U N
-Quantum Mechanics: Concepts and Applications N. Zettili 2001 [33] U Y
-Elements of quantum mechanics M. Fayer 2001 [34] U N
-Understanding quantum mechanics R. Omnés 1999 [35] U N
-Particles behave like waves T. Moore 1998 [36] U N
-Introduction to quantum mechanics D. Griffiths 1995 [37] U Y
-Quantum theory A. Peres 1993 [38] G Y
-Quantum mechanics F. Mandl 1992 [39] U N
-A modern approach to quantum mechanics J. Townsend 1992 [40] U Y
-Quantum mechanics: an introduction W. Greiner 1989 [41] U N
-Quantum mechanics A. Rae 1986 [15] U N
-Modern quantum mechanics J. Sakurai 1985 [42] G Y
-Principles of Quantum mechanics R. Shankar 1980 [43] G Y
-An introduction to quantum physics A. French & E. Taylor 1978 [44] U Y
-Quantum mechanics C. Cohen-Tannoudji 1977 [45] G Y
-Quantum physics of atoms, molecules, solids, nuclei and particles R. Eisberg & R. Resnick 1974 [46] U Y
-Quantum physics S. Gasiorowicz 1974 [47] U Y
-Feynman lectures on physics volume 3 R. Feynman 1965 [48] U Y
-Quantum mechanics A. Davydov 1965 [49] G Y
-Elementary quantum mechanics P. Fong 1962 [50] U N
-Quantum mechanics E. Merzbacher 1961 [51] G N
-Introduction to quantum mechanics R. Dicke & J. Wittke 1960 [52] G N
-Quantum mechanics A. Messiah 1959 [53] G N
-Quantum mechanics: non-relativistic theory L. Landau & L. Lifshitz 1958 [54] G Y
-Mathematical foundations of quantum mechanics J. von Neumann 1955 [55] G Y
-Quantum theory D. Bohm 1951 [56] G N
-Quantum mechanics L. Schiff 1949 [57] G N
-The principles of quantum mechanics P. Dirac 1930 [58] G Y
-Quantum mechanics: a modern development L. Ballentine 1998 [14] G N
-Quantum computation and quantum information M. Nielsen & I. Chuang 2000 [18] G N
-An introduction to quantum theory K. Hannabuss 1997 [17] G N
-Conceptual foundations of quantum physics D. Home 1997 [6] - N
+Lecture notes on quantum mechanics D. Tong 2025 [tong2025] U N
+Basic quantum mechanics K. Tamvakis 2019 [tamvakis2019] U N
+A first introduction to quantum physics P. Kok 2018 [kok2023] U N
+Lectures on quantum mechanics S. Weinberg 2013 [weinberg2013] G N
+The physics of quantum mechanics J. Binney & D. Skinner 2013 [binney2013physics] U Y
+Quantum concepts in physics M. Longair 2013 [longair_2013] G N
+Quantum mechanics D. McIntyre et al 2012 [mcintyre2012quantum] U Y
+Quantum mechanics: theory and experiment M. Beck 2012 [beck2012] U N
+Quantum mechanics: a new introduction K. Konishi & G. Paffuti 2009 [Konishi:2009qva] U N
+Quantum mechanics for scientists and engineers D. Miller 2008 [miller2010] U N
+Quantum mechanics: a textbook for undergraduates M. Jain 2007 [jain2017] U N
+Introduction to quantum mechanics H. Müller-Kirsten 2006 [muller-kisten2006] U Y
+Quantum physics M. Le Bellac 2006 [bellac2011quantum] G N
+Lectures on quantum mechanics J. Basdevant 2005 [basdevant2023] G N
+Introduction to quantum mechanics A. Philips 2003 [phillips2013introduction] U N
+Quantum Mechanics: Concepts and Applications N. Zettili 2001 [zettili2001] U Y
+Elements of quantum mechanics M. Fayer 2001 [fayer2001] U N
+Understanding quantum mechanics R. Omnés 1999 [Omnes1999-OMNUQM] U N
+Particles behave like waves T. Moore 1998 [moore1998] U N
+Introduction to quantum mechanics D. Griffiths 1995 [Griffiths2004Introduction] U Y
+Quantum theory A. Peres 1993 [peres1993] G Y
+Quantum mechanics F. Mandl 1992 [mandl2013quantum] U N
+A modern approach to quantum mechanics J. Townsend 1992 [townsend2012modern] U Y
+Quantum mechanics: an introduction W. Greiner 1989 [greiner1989] U N
+Quantum mechanics A. Rae 1986 [rae2002] U N
+Modern quantum mechanics J. Sakurai 1985 [Sakurai:1167961] G Y
+Principles of Quantum mechanics R. Shankar 1980 [Shankar:102017] G Y
+An introduction to quantum physics A. French & E. Taylor 1978 [french1978] U Y
+Quantum mechanics C. Cohen-Tannoudji 1977 [Cohen-Tannoudji:101367] G Y
+Quantum physics of atoms, molecules, solids, nuclei R. Eisberg & R. Resnick 1974 [eisberg1974] U Y
+and particles
+Quantum physics S. Gasiorowicz 1974 [gasiorowicz2007quantum] U Y
+Feynman lectures on physics volume 3 R. Feynman 1965 [feynman1977feynman] U Y
+Quantum mechanics A. Davydov 1965 [davydov1965] G Y
+Elementary quantum mechanics P. Fong 1962 [fong2013elementary] U N
+Quantum mechanics E. Merzbacher 1961 [merzbacher1998] G N
+Introduction to quantum mechanics R. Dicke & J. Wittke 1960 [dicke1960] G N
+Quantum mechanics A. Messiah 1959 [messiah1999quantum] G N
+Quantum mechanics: non-relativistic theory L. Landau & L. Lifshitz 1958 [Landau1981Quantum] G Y
+Mathematical foundations of quantum mechanics J. von Neumann 1955 [vonneumann1955] G Y
+Quantum theory D. Bohm 1951 [bohm1951] G N
+Quantum mechanics L. Schiff 1949 [schiff1955] G N
+The principles of quantum mechanics P. Dirac 1930 [Dirac1930-DIRTPO] G Y
+Quantum mechanics: a modern development L. Ballentine 1998 [ballentine1998quantum] G N
+Quantum computation and quantum information M. Nielsen & I. Chuang 2000 [nielsen2010] G N
+An introduction to quantum theory K. Hannabuss 1997 [Hannabuss:1999fa] G N
+Conceptual foundations of quantum physics D. Home 1997 [Home1997] - N
 4 Analysis from example texts
 To illustrate the analysis process we will consider several texts as examples.
-These are chosen because they all share a common case of having to infer agreement with at least one proposition (usually “measurement creates observed values”).
 This allows the reader to determine that we are carefully evidencing degrees of assent or dissent in each case.
@@
 4.1 Binney & Skinner
-We begin with The physics of quantum mechanics by Binney & Skinner [22].
+We begin with The physics of quantum mechanics by Binney & Skinner [binney2013physics].
 For “measurement disturbs systems” we find initial evidence on page 3: “Moreover, the act of measurement inevitably disturbs the atom, and leaves it in a different state from the one it was in before we made the measurement” (note that this is not in the context of the physics of a particular type of measurement).
@@
 For “probability from uncertainty” the authors dissent, as we have seen that they attribute probability to the uncertain state of the measuring apparatus and its subsequent disturbing of the system.
-For “measurement creates observed values” we infer assent, their view seems to be that systems are disturbed by measurement, changing the values found rather than creating them altogether.
+For “no pre-measurement values”, their view on pages 130–131 seems to be that systems are disturbed by measurement, changing the values found rather than creating them altogether (the Born rule is stated as a probability of finding a value).
 We have already seen that the authors view the state as ontologically complete, an “extended” theory is possible, but it merely accounts for the actual physics of measurement (page 130 and 131).
+There is no mention of other viable interpretations.
 4.2 Tamvakis
-Now we consider Basic quantum mechanics by Tamvakis [19].
+Now we consider Basic quantum mechanics by Tamvakis [tamvakis2019].
 On page 86 it is outlined, as part of the axioms of quantum mechanics, that the state changes during measurement indicating support for “measurement disturbs systems”.
@@
 Page 5 describes the uncertainty principle in terms of knowledge of $x$ and $p_{x}$ “thus, viewed from a classical viewpoint, in a microscopic system, complete knowledge of $x$ and $p_{x}$ is impossible”, we infer this is indicating that probabilities result from incomplete knowledge implied by the uncertainty principle.
-For “measurement creates observed values”, we note that it is stated on page 86 that a superposition represents the pre-measurement state, since this cannot have a definite property value we must conclude that definite property values are created, or at least changed, during measurement.
-This will satisfy an assent to “measurement observed creates values” as a pre-measurement state with definite values cannot be recovered, as the author describes measurement as “acausal” on page 87.
+For “no pre-measurement values”, we note that it is stated on page 86 that a superposition represents the pre-measurement state, since this cannot have a definite property value we must conclude that definite property values are created, or at least changed, during measurement.
+This will satisfy an assent to the proposition as the author describes measurement as “acausal” on page 87, so the initial state is irrelevant (the Born rule is stated as a probability of finding a value).
+There is no mention of other viable interpretations.
 4.3 Zettili
-Next we consider Quantum Mechanics: Concepts and Applications by Zettili [33].
+Next we consider Quantum Mechanics: Concepts and Applications by Zettili [zettili2001].
 On page 24, “measurement interferes with the state of microscopic objects” and “it is the effects of interference by the equipment on the system which is the essence of quantum mechanics” on page 165 demonstrate support for measurement disturbance.
@@
 We combine previous statements from 165 and 169 to conclude that probability is introduced by measurement disturbance (uncertainty).
-Based on the axioms from 158 we infer that measurement creates observed values, as pre-measurement states can be superpositions and apply to individual particles/realisations.
+Based on the axioms from 158, where the Born rule is stated as a probability of finding a value, as well as the state being changed to an eigenstate on measurement, we see assent to “no pre-measurement values”.
+There is no mention of other viable interpretations.
 4.4 McIntyre, Manogue, & Tate
-We now turn to Quantum mechanics by McIntyre, Manogue, and Tate [24].
+We now turn to Quantum mechanics by McIntyre, Manogue, and Tate [mcintyre2012quantum].
 On page 8 we find that measurements disturbing systems is a “key feature” of quantum mechanics.
@@
 The same discussion refers the reader to Feynman and Cohen-Tannoudji which seems to indicate the ultimate outcome is individual particles are wave-like too.
-Page 192 indicates we cannot know a particles trajectory but does not deny its existence.
+Page 192 indicates we cannot know a particle’s trajectory but does not deny its existence.
 Page 48 includes collapse in the postulates.
@@
 “The state of quantum mechanical system, including all the information you can know about it, is represented mathematically by a normalised ket $\left|\psi\right$>” covers ontological completeness.
-Discussion of Schrödinger’s cat on 167 and 168 is illuminating in terms of the role of ensembles.
-Here it is asserted that only decoherence prevents superpositions at classical scales and interference is used to ascribe physical significance to superposition, beyond mere statistics.
-Thus, we see that the authors take superpositions to describe the state of individual realisations.
-Acceptance of real superpositions combined with the stated quantum postulates requires observed values are created by measurement.
+On page 14 the authors describe the Born rule as yielding the probability to measure a given value.
+On page 19, they say superpositions are not probabilistic mixtures.
+These last two, combined with page 8’s description of superposed states as having indefinite values indicates assent to “no pre-measurement values”.
+The authors indicate that matters of interpretation are irrelevant, and advise the reader to “shut up and calculate” on page 168.
 4.5 Konishi & Paffuti
-Finally we review Quantum mechanics: a new introduction by Konishi & Paffuti [26].
+Finally, we review Quantum mechanics: a new introduction by Konishi & Paffuti [Konishi:2009qva].
 On page 29 the authors support the notion that all measurement causes disturbance by stating that “the disturbance caused by the process of measurement on the microscopic system under examination cannot be controlled beyond a certain limit.” The unqualified illustration of the uncertainty principle via “Heisenberg microscopes” on page 28 indicates the authors link uncertainty to measurement disturbance.
@@
 These indicate that quantum states do not apply to ensembles but to single realisations of systems.
-On page 29 the authors agree that probability derives from the uncertainty principle “the probabilistic nature of quantum mechanical prediction for the results of measurement is an inevitable consequence of Heisenberg’s uncertainty principle.” Although the discussion on 518 and 519 is ambiguous to whether values are being revealed or created (the authors use “indefinite” properties of a superposition here to mean “unpredictable”), we must infer that they are at least being altered by measurement to comport with previous assertions.
+On page 29 the authors agree that probability derives from the uncertainty principle “the probabilistic nature of quantum mechanical prediction for the results of measurement is an inevitable consequence of Heisenberg’s uncertainty principle.” Although the discussion on 518 and 519 is ambiguous to whether values are being revealed or created (the authors use “indefinite” properties of a superposition here to mean “unpredictable”), we must infer that there are no pre-measurement values following the assertion, on page 23, that superposition properties are intermediate to definite states.
+This is reinforced by phrasing the Born rule in terms of “finding property values” on page 31.
 On page 21 the authors state that “knowledge of the wave function amounts to complete knowledge of the state” which, combined with prior assertions fully supports ontological completeness.
+There is no mention of other viable interpretations.
 5 Results and discussion
-In Table 2 we display all the data used for this study.
+In Table 2 we display all the data used for this study (pages referenced are listed in Table 3).
 Each work’s response to the propositions is characterised by a degree of agreement.
@@
 […]
-Tong [16] 1 1 1 1 1 1 -1 0 1 1 0.5 1 2,3,4,5,13,14,16,17,18,21,78
-Tamvakis [19] 1 0 1 1 0 1 1 1 1 1 0.5 1 5,6,7,86,87,442
-Kok [20] 1 1 1 1 1 1 1 0 1 0 1 0 6,182,186,211,58,206,225,226
-Weinberg [21] 0 0 0.5 -1 0 -1 0 0 1 0 0 -1 21,25,82,83,88,95,97
-Binney & Skinner [22] 1 1 1 -1 0 -1 -1 0 1 -1 0.5 1 3,7,8,14,15,28,71,130,131
-Longair [23] 1 1 1 1 0.5 1 1 1 1 1 0.5 1 178,183,359,360,363,367
-Mcintyre et al [24] 1 0.5 1 1 -1 1 1 1 1 0 0.5 1 7,8,48,59,121,167,168,192,207,208
-Beck [25] 1 0 1 1 0 1 0 1 1 0 1 1 62,91,92,101,129
-Konishi & Paffuti [26] 1 1 1 1 1 1 1 0 1 1 0.5 1 6,21,23,25,26,28,29,518,519
-Miller [27] 1 0 1 1 0 1 1 0 1 0 1 1 6,9,74,83,84,425,450
-Jain [28] 1 1 1 1 0 1 1 1 1 0 1 1 61,77,83,86,89,209,210,214
-Müller-Kirsten [29] 1 0 1 1 0 0 1 1 1 1 1 1 9,13,17,19,20,169
-Le Bellac [30] 1 0 1 1 0 1 1 0 1 0 1 1 22,24,97,101,102,104,179,185
-Basdevant [31] 1 -1 1 1 1 1 1 1 1 0 1 1 3,17,20,35,43,47,73,79,171,420
-Phillips [32] 1 1 1 1 1 1 1 0 1 0 1 -1 6,11,13,16,21,27,50
-Zettili [33] 1 1 1 1 -0.5 1 1 1 1 0.5 0.5 1 18,24,25,26,28,30,51,158,165,169
-Fayer [34] 1 1 1 1 1 1 1 0 1 0 1 1 1,2,3,5,32,33,42,55
-Omnés [35] 1 1 1 1 1 1 1 1 1 1 1 1 31,45,47,49,53,91
-Moore [36] 1 0 -1 1 0.5 1 1 1 1 0 1 -1 69,77,81,106,107,108,135
-Griffiths [37] 1 1 -1 0 0.5 1 1 0 1 -1 1 1 1,2,3,4,5,19,112,420,431
-Peres [38] -1 0 -1 1 -1 0 1 1 -1 -1 1 1 5,11,14,20,24,25,26,27,93
-Townsend [40] 1 1 1 1 0 1 0 0 1 0 1 1 6,8,10,13,16,25,165,211
-Mandl [39] 1 1 0.5 1 0 0 0 0 0.5 0 0 1 3,50,83
-Greiner [41] 1 0 1 1 1 1 1 0.5 1 0 1 1 29,34,38,51,53,469,473
-Rae [15] 0.5 -1 -1 1 1 1 1 1 1 0 1 1 6,8,10,61,73,79,261,273,287
-Sakurai [42] 1 1 -1 -1 0 1 0 0 1 0 0 1 6,11,23,24,101
-Shankar [43] 1 1 1 1 1 1 1 0 1 0 0.5 1 111,112,113,115,116,129,140
-French [44] 1 1 1 1 0.5 1 1 0 1 0 0.5 0 72,76,81,85,92,93,243,317,328
-Cohen-Tannoudji [45] 1 1 1 1 1 1 1 1 1 0 0.5 1 14,19,21,27,28,221,252,253,279
-Eisberg & Resnick [46] 1 1 1 1 0 0 1 1 1 1 0.5 0 64,65,68,70,73,74,76,87,88,147
-Gasiorowicz [47] 1 1 1 1 0 0.5 1 0 1 0 1 1 13,24,32,94,318,323,327
-Feynman [48] 1 1 1 1 1 1 1 1 1 1 1 1 1-(6,7,9,10,11),2-(1,2,3,8,10)
-Davydov [49] 1 0 1 1 1 1 0 1 1 0 1 1 2,3,4,6,33,37,40,51
-Fong [50] 1 1 1 1 0 1 1 1 1 1 1 1 28,32,50,51,52,53,114,116,123,124
-Merzbacher [51] 0 0 1 1 1 -1 1 1 1 0 1 1 3,4,6,9,17,18,28,408
-Dicke & Wittke [52] 1 1 1 1 0 1 1 1 1 0 1 1 14,15,16,24,27,40,74,91,110,115,120
-Messiah [53] 1 1 1 1 1 1 1 1 1 0 0.5 1 20,21,49,116,119,140,143,144,153,157,162
-Landau & Lifshitz [54] 1 1 1 1 1 1 0 0 1 1 1 1 2,3,4,5,6,24,47,157
-von Neumann [55] 1 0 1 0 1 1 -1 0 1 1 1 1 4,210,214,216,247,326,403
-Bohm [56] 1 1 1 1 1 1 1 1 1 1 0.5 1 59,73,80,99,104,117,120,145,160
-Schiff [57] 1 1 1 1 1 0 1 1 1 0 0.5 1 3,6,7,8,13,16,21
-Dirac [58] 1 1 1 1 0 1 1 0 1 0 1 1 4,7,8,9,16,34,46,52,98
-In Figure 1, we see the agreement frequencies for each of the questions in turn.
-We display these in terms of 4 samples: the full sample (black), undergraduate books (yellow), graduate texts (blue), works on the ‘Notable books’ list from Wikipedia (red), pre-1980 works (gray), post-1980 works (green), and post-2000 works (magenta).
-The only question substantially impacted by inference of response is “measurement” creates values.
-We supplement this plot with Fig.
-2, which shows that rejection is not inferred to any significant degree.
+Tong 2025 1 1 1 1 1 1 -1 0 1 1 1 1 1
+Tamvakis 2019 1 0 1 1 0 1 1 1 1 1 1 1 1
+Kok 2018 1 1 1 1 1 1 1 0 1 0 0 1 0.5
+Weinberg 2013 0 0 0.5 -1 0 -1 0 0 1 0 -1 1 -1
+Binney and Skinner 2013 1 1 1 -1 0 -1 -1 0 1 -1 1 -1 1
+Longair 2013 1 1 1 1 0.5 1 1 1 1 1 1 1 1
+McIntyre, Manogue, and Tate 2012 1 0.5 1 1 -1 1 1 1 1 0 1 1 1
+Beck 2012 1 0 1 1 0 1 0 1 1 0 1 1 1
+Konishi and Pafutti 2009 1 1 1 1 1 1 1 0 1 1 1 1 1
+Miller 2008 1 0 1 1 0 1 1 0 1 0 1 1 -1
+Jain 2007 1 1 1 1 0 1 1 1 1 0 1 1 1
+Muller-Kirsten 2006 1 0 1 1 0 0 1 1 1 1 1 1 1
+Le Bellac 2006 1 0 1 1 0 1 1 0 1 0 1 1 1
+Basdevant 2005 1 -1 1 1 1 1 1 1 1 0 1 1 1
+Phillips 2003 1 1 1 1 1 1 1 0 1 0 -1 1 1
+Zettili 2001 1 1 1 1 -0.5 1 1 1 1 0.5 1 1 1
+Fayer 2001 1 1 1 1 1 1 1 0 1 0 1 1 1
+Omnes 1999 1 1 1 1 1 1 1 1 1 1 1 1 -1
+Moore 1998 1 0 -1 1 0.5 1 1 1 1 0 -1 1 1
+Griffiths 1995 1 1 -1 0 0.5 1 1 0 1 -1 1 1 0.5
+Peres 1993 -1 0 -1 1 -1 0 1 1 -1 -1 1 1 0.5
+Mandl 1992 1 1 0.5 1 0 0 0 0 0.5 0 1 -1 1
+Townsend 1992 1 1 1 1 0 1 0 0 1 0 1 1 1
+Greiner 1989 1 0 1 1 1 1 1 0.5 1 0 1 -1 -1
+Rae 1986 0.5 -1 -1 1 1 1 1 1 1 0 1 1 -1
+Sakurai 1985 1 1 -1 -1 0 1 0 0 1 0 1 1 0.5
+Shankar 1980 1 1 1 1 1 1 1 0 1 0 1 1 1
+French and Taylor 1978 1 1 1 1 0.5 1 1 0 1 0 0 1 1
+Cohen-Tannoudji 1977 1 1 1 1 1 1 1 1 1 0 1 -1 1
+Eisberg and Resnick 1974 1 1 1 1 0 0 1 1 1 1 0 1 1
+Gasiorowicz 1974 1 1 1 1 0 0.5 1 0 1 0 1 1 1
+Feynman 1965 1 1 1 1 1 1 1 1 1 1 1 1 1
+Davydov 1965 1 0 1 1 1 1 0 1 1 0 1 1 1
+Fong 1962 1 1 1 1 0 1 1 1 1 1 1 1 -1
+Merzbacher 1961 0 0 1 1 1 -1 1 1 1 0 1 1 1
+Dicke and Wittke 1960 1 1 1 1 0 1 1 1 1 0 1 1 1
+Messiah 1959 1 1 1 1 1 1 1 1 1 0 1 1 1
+Landau and Lifshitz 1958 1 1 1 1 1 1 0 0 1 1 1 1 1
+von Neumann 1955 1 0 1 0 1 1 -1 0 1 1 1 1 1
+Bohm 1951 1 1 1 1 1 1 1 1 1 1 1 1 1
+Schiff 1949 1 1 1 1 1 0 1 1 1 0 1 1 1
+Dirac 1930 1 1 1 1 0 1 1 0 1 0 1 1 1
+Ballentine 1998 -1 -1 -1 -1 -1 -1 -1 0 -1 -1 -1 -1 -1
+Nielsen and Chuang 2000 1 0 -1 0 0 1 0 0 1 0 1 1 1
+Hannabuss 1997 1 0 0.5 1 0 1 1 0 1 0 -1 -1 1
+Home 1997 1 0 -1 1 1 0 1 0 1 0 1 1 -1
+Table 3: Table of the pages used in studied references.
+Author Year Pages
+Tong 2025 10,11,20,21,25,29,97,100,108,123,124,574
+Tamvakis 2019 5,6,7,86,87,442
+Kok 2018 6,58,182,186,206,211,225,226,264
+Weinberg 2013 21,25,26,82,83,88,95,97
+Binney and Skinner 2013 3,7,8,14,15,28,71,130,131
+Longair 2013 178,183,346,359,360,363,367
+McIntyre, Manogue, and Tate 2012 7,8,14,19,48,59,121,167,168,192,207,208
+Beck 2012 62,91,92,93,101,129,218
+Konishi and Pafutti 2009 6,21,23,25,26,28,29,518,519
+Miller 2008 6,9,74,83,84,425,449,450
+Jain 2007 61,77,83,86,89,209,210,214
+Muller-Kirsten 2006 9,13,17,19,20,77,169
+Le Bellac 2006 22,24,97,101,102,104,179,185
+Basdevant 2005 3,17,20,35,43,47,73,79,171,420
+Phillips 2003 6,11,13,16,21,27,50
+Zettili 2001 18,24,25,26,28,30,51,158,165,169
+Fayer 2001 1,2,3,5,32,33,42,55
+Omnes 1999 31,45,47,49,53,91
+Moore 1998 69,77,81,106,107,108,113,135
+Griffiths 1995 1,2,3,4,5,19,112,420,431
+Peres 1993 5,11,14,20,24,25,26,27,93
+Mandl 1992 3,13,50,83
+Townsend 1992 6,8,10,13,16,25,165,211
+Greiner 1989 29,34,38,51,53,259,469,473
+Rae 1986 6,8,10,61,73,79,261,273,287
+Sakurai 1985 6,11,23,24,101
+Shankar 1980 111,112,113,115,116,129,140
+French and Taylor 1978 72,76,81,85,92,93,120,243,317,328
+Cohen-Tannoudji 1977 14,19,21,27,28,221,252,253,279
+Eisberg and Resnick 1974 64,65,68,70,73,74,76,87,88,147
+Gasiorowicz 1974 13,24,28,32,94,318,323,327,328
+Feynman 1965 1-10,1-11,1-6,1-7,1-9,2-1,2-10,2-2,2-3,2-8
+Davydov 1965 2,3,4,6,33,37,40,51
+Fong 1962 28,29,32,50,51,52,53,114,116,123,124
+Merzbacher 1961 3,4,6,9,17,18,28,408
+Dicke and Wittke 1960 14,15,16,24,25,27,40,74,91,110,115,120
+Messiah 1959 20,21,46,49,116,117,119,140,143,144,153,157,162
+Landau and Lifshitz 1958 2,3,4,5,6,24,31,47,57,157
+von Neumann 1955 4,210,214,216,247,326,403
+Bohm 1951 59,73,80,92,99,104,117,120,145,160
+Schiff 1949 3,6,7,8,13,16,21
+Dirac 1930 4,7,8,9,16,34,46,52,98
+Ballentine 1998 100,133-136,138,226,227,233,236,244,247,45,47,55
+Nielsen and Chuang 2000 13,15,80,81,88,89,96
+Hannabuss 1997 3,11,77,86,173,177
+Home 1997 14,16,18,19,22,38,52
+In Figure 1 we display the agreement frequencies as a function of time via 3 subsamples: pre-1980 works (gray), post-1980 works (green), and post-2000 works (magenta).
+We also display the following 4 samples in Fig.
+2: the full sample (black), undergraduate books (yellow), graduate texts (blue), and works on the ‘Notable books’ list from Wikipedia (red).
+We supplement this with rejection frequencies in Figs.
+3 and 4.
 It would be ideal to include the significance of the consensus in each case.
@@
 Given a sample of $42$ works, this is sign of substantial consensus.
-Figure 1: Affirmation frequency for responses to each question.
-The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), pre-1980 (grey), post-1980 (green), and post-2000 (magenta).
+Figure 1: Affirmation frequency for responses to each question as a function of time.
+The bars represent different samples: pre-1980 (grey), post-1980 (green), and post-2000 (magenta).
 The error bars reflect the effect of including values inferred for logical consistency.
-Figure 2: Rejection frequency for responses to each question.
-The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red), pre-1980 (grey), post-1980 (green), and post-2000 (magenta).
+Figure 2: Affirmation frequency for responses to each question for various subsamples.
+The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), Wikipedia notable (red).
+The error bars reflect the effect of including values inferred for logical consistency.
+Figure 3: Rejection frequency for responses to each question as a function of time.
+The bars represent different samples: pre-1980 (grey), post-1980 (green), and post-2000 (magenta).
+The error bars reflect the effect of including values inferred for logical consistency.
+Figure 4: Rejection frequency for responses to each question for various subsamples.
+The bars represent different samples: all (black), undergraduate (yellow), graduate (blue), and Wikipedia notable (red).
 The error bars reflect the effect of including values inferred for logical consistency.
@@
 Interestingly, we see that not all the books regard $\psi$ as ontologically complete.
-However, the support for this proposition is robust with assent frequency $0.85$ in the full sample, with all subsamples having assent frequency $\geq 0.8$.
+However, the support for this proposition is robust with assent frequency $\sim 0.85$ in the full sample, with all subsamples having assent frequency $\geq 0.8$.
 It is interesting that the lowest support for this occurs within the undergraduate and post-2000 samples, with outright rejection passing $0.1$ for post-1980 and post-2000 subsamples.
 Notably, there is no rejection in the pre-1980 sample, giving an indication that responses to this proposition are evolving.
-In terms of dissent, Phillips [32] and Moore [36] limit themselves to saying that $\psi$ can determine all the probabilities associated with a system but do not go on to say this is all the information allowed.
-Whereas, [21] openly suggests that interpretive difficulties might be overcome by finding a more general theory to supersede quantum mechanics.
+In terms of dissent, Phillips [phillips2013introduction] and Moore [moore1998] limit themselves to saying that $\psi$ can determine all the probabilities associated with a system but do not go on to say this is all the information allowed.
+Whereas, [weinberg2013] openly suggests that interpretive difficulties might be overcome by finding a more general theory to supersede quantum mechanics.
 We also see strong support (a minimum frequency of $0.94$) for the notion that a quantum state describes a single instance (as opposed to an ensemble).
 Notably, this achieves $0.96$ assent in the full sample and the undergraduate works are near unanimous on this count.
+Only Peres [peres1993] dissents from the notion that quantum states describe individual instances.
 In state collapse we find an interesting case.
@@
 The collapse proposition still has subsample assent frequencies above $0.75$ while achieving $0.79$ in the full sample.
+The rate of rejection for collapse increases substantially when the post-1980 (and post-2000) subsample is compared to the pre-1980 case.
+However, the assent frequency for collapse also increases from pre to post-1980, perhaps a marker of increased controversy.
+It is also noteworthy that, in a sample of $42$ works, there are only three dissenters to the notion that quantum states collapse: Weinberg [weinberg2013], Merzbacher [merzbacher1998], as well as Binney & Skinner [binney2013physics].
+Additionally, the topic is not mentioned by Peres, Mandl, Müller-Kirsten, and Schiff [peres1993, mandl2013quantum, muller-kisten2006, schiff1955].
+Interestingly, we note that Home [Home1997] argues that collapse is not essential to characterise the orthodoxy.
+However, this is strongly disputed by the data gathered here.
 From these results we can also see that an orthodox view requires the state to contain the entire ontological situation and apply to individual particles.
@@
 There are tentative signs of evolution in the consensus here.
-Particularly, the rate of rejection for collapse and ontological completeness both increase substantially when the post-1980 (and post-2000) subsample is compared to the full set.
-However, the assent frequency for collapse also increases from pre to post-1980, perhaps a marker of increased controversy.
-It is also noteworthy that, in a sample of $42$ works, there are only three dissenters to the notion that quantum states collapse: Weinberg [21], Merzbacher [51], as well as Binney & Skinner [22].
-Additionally, the topic is not mentioned by Peres, Mandl, Müller-Kirsten, and Schiff [38, 39, 29, 57].
-Only Peres [38] dissents from the notion that quantum states describe individual instances.
-Interestingly, we note that Home [6] argues that collapse is not essential to characterise the orthodoxy.
-However, this is strongly disputed by the data gathered here.
-Overall, we see that the orthodoxy has it that the quantum state is ontologically complete, refers to individual instances of systems, and collapses on measurement.
 Now we turn to propositions on the nature of particles.
 A strong degree of consensus is clear around the notion that individual particles display wave-like behaviour with assent frequency $0.88$ in the full sample, while maintaining subsample frequencies $\geq 0.78$.
-The only dissenters here being Weinberg [21], Binney & Skinner [22], and Sakurai [42].
-The topic is also not discussed in von Neumann [55] (perhaps unsurprising given the mathematical focus) and Griffiths [37].
+The only dissenters here being Weinberg [weinberg2013], Binney & Skinner [binney2013physics], and Sakurai [Sakurai:1167961].
+The topic is also not discussed in von Neumann [vonneumann1955] (perhaps unsurprising given the mathematical focus) and Griffiths [Griffiths2004Introduction].
 Interestingly, the lowered assent in the post-1980 subsample seems to stem from more recent graduate texts rejecting the proposition.
@@
 Reasonable support also exists for the notion of wave-particle duality with assent $0.78$ in the full sample and subsample frequency $\geq 0.71$.
-Here only von Neumann [55], Binney & Skinner [22], and Tong [16] outright deny this proposition.
-It is also not mentioned by Weinberg [21], Mandl [39], Townsend [40], Davydov [49] and Sakurai [42].
+Here only von Neumann [vonneumann1955], Binney & Skinner [binney2013physics], and Tong [tong2025] outright deny this proposition.
+It is also not mentioned by Weinberg [weinberg2013], Mandl [mandl2013quantum], Townsend [townsend2012modern], Davydov [davydov1965] and Sakurai [Sakurai:1167961].
 There is discernable evolution based on publication year, rejection is much more common in the post-2000 subsample than either post-1980 or pre-1980 cases.
@@
 However, it is not rejected at all in the pre-1980 or post-2000 cases.
-The works between 1980 and 2000 that cause this are Griffiths [37], Peres [38], McIntyre [24], Moore [36], and Sakurai [42].
+The works between 1980 and 2000 that cause this are Griffiths [Griffiths2004Introduction], Peres [peres1993], McIntyre [mcintyre2012quantum], Moore [moore1998], and Sakurai [Sakurai:1167961].
 This shows some evidence of a change in the consensus over time, but it appears to have been a fluctuation, returning towards the pre-1980 consensus after two decades.
-Home [6] does not include this proposition in their conception of the orthodoxy, likely as it is a flawed proposition.
-As argued in [9], the uncertainty principle could not function in the suggested manner.
+Home [Home1997] does not include this proposition in their conception of the orthodoxy, likely as it is a flawed proposition.
+As argued in [Popper1982-POPQTA], the uncertainty principle could not function in the suggested manner.
 If the measurements from individual particles had the uncertainty principle’s scatter, the final ensemble distribution would have a larger uncertainty.
@@
 The idea is never rejected pre-1980, but rejection appears afterwards and is largely maintained in more recent works.
-Asserting there are no trajectories still finds agreement in modern works like Kok [20], Miller [27], Jain [28], Basdevant [31], Konishi & Paffuti [26], Cohen-Tannoudji [45], Shankar [43], Greiner [41], and Phillips [32].
-There is also partial agreement in Moore [36] and Longair [23].
+Asserting there are no trajectories still finds agreement in modern works like Kok [kok2023], Miller [miller2010], Jain [jain2017], Basdevant [basdevant2023], Konishi & Paffuti [Konishi:2009qva], Cohen-Tannoudji [Cohen-Tannoudji:101367], Shankar [Shankar:102017], Greiner [greiner1989], and Phillips [phillips2013introduction].
+There is also partial agreement in Moore [moore1998] and Longair [longair_2013].
 However, the mode response is 0 in the undergraduate, ‘Wikipedia notable’, and post-1980 samples.
@@
 Thus, we see that the orthodoxy holds that particles are individually wave-like, have a dual wave-particle nature reflected in their statistics, and that the uncertainty principle applies to measurements performed on single instances of systems (rather than acting only as a statistical deviation over an ensemble).
-Next, we turn to propositions around the question of why quantum mechanics predicts only statistical results.
+Next, we turn to propositions around the question of why quantum mechanics predicts only probabilistic results.
 There is clear consensus, regardless of whether we confine our attention to modern works, that measurement disturbs quantum systems with assent $0.91$ in the full sample, with a minimum of $0.81$ in the graduate case.
 Notably, the pre-1980 and post-2000 samples have very similar assent frequencies.
-Indeed, only Peres [38] actively dissents.
-Their dissent is justified by the possibility of non-interactive measurement in QM [59].
-There is also support for the idea that measurement itself actually creates the observed values with frequency $0.76$ (without inferred values this drops to $0.59$) in the full sample while retaining 0.730.73($0.58$) in the post-1980 subsample.
-Interestingly, the assent frequency drops further in the post-2000 case, indicating that there is some evolution on this proposition.
-However, it is never rejected in any sample.
-It should be noted that the affirmation is only stated openly around 6060% of the time, despite being a plausible corollary to “measurement disturbs quantum systems” and that there are no active dissenters, not even [38].Only three authors do not provide grounds for inferring agreement: Sakurai [42], Mandl [39], and Weinberg [21].
-The common case for inference is as follows: $\psi$ applies to individual systems, hence superpositions are real.
-Measurement disturbs the system and changes a superposition to an eigenstate, so it is either changing an unknown underlying value or creating a definite one where there was none.
-In either case the pre-existing value is totally superfluous if the collapse process is irreversible and indeterminate.
-However, the statistics here are significantly affected by the reliability of the inferences being made.
+Indeed, only Peres [peres1993] actively dissents.
+Their dissent is justified by the possibility of non-interactive measurement in QM [Elitzur_1993].
+There is also support for the idea that quantum systems have no pre-measurement properties, with assent $\sim 0.88$ in the full sample.
+For time evolution, this has >0.9>0.9 support pre-1980 but retains >0.8>0.8 in later samples.
+This proposition is most rejected in the post-1980 sample, at $\approx 0.15$, dropping to $\approx 0.12$ in the post-2000 sample.
+This is still twice as common as the pre-1980 works.
+It is rejected by Binney & Skinner [binney2013physics], Mandl [mandl2013quantum], and Cohen-Tannoudji [Cohen-Tannoudji:101367].
+Of these, Mandl and Cohen-Tannoudji describe the Born rule as $|\psi(x)|^{2}$ giving the probability for a particle being at $x$, instead of merely “found” there.
 The notion that probabilities in QM result from the uncertainty principle has always been a minority view it appears, as most authors do not mention anything like it, achieving a maximum assent of $0.4$ in the pre-1980 sample.
 However, it is not rejected until after 1980 with the rejection rate dropping substantially in the post-2000 works.
-This coincides with an increase in assent post-2000, where Tong [16], Tamvakis [19], Longair [23], Konishi & Paffuti [26], and Müller-Kirsten [29] all support the position.
-In the same sample, only Binney & Skinner [22] reject it, so the mode response is still to leave it out.
+This coincides with an increase in assent post-2000, where Tong [tong2025], Tamvakis [tamvakis2019], Longair [longair_2013], Konishi & Paffuti [Konishi:2009qva], and Müller-Kirsten [muller-kisten2006] all support the position.
+In the same sample, only Binney & Skinner [binney2013physics] reject it, so the mode response is still to leave it out.
 There is still some support for the notion that the uncertainty principle is a consequence of measurement disturbance, with assent frequency $0.69$.
@@
 Interestingly, the sample with most affirmation for this last proposition is the “Wikipedia notable” one, with $0.79$ assent.
-The only outright rejections here are Rae [15] and Basdevant [31], who both note the uncertainties can happily be measured in separate runs of an experiment, so “disturbance” need not occur.
+The only outright rejections here are Rae [rae2002] and Basdevant [basdevant2023], who both note the uncertainties can happily be measured in separate runs of an experiment, so “disturbance” need not occur.
 Thus, for the orthodoxy, we do not reveal the values of studied quantum systems but force them to assume definite values (or alter them) via some uncontrolled disturbance attached to the abstract notion of measurement.
-The “uncontrolled” aspect of this is vital to square the statistical nature of QM with the proposition that quantum states apply to individual instances of systems.
+The “uncontrolled” aspect of this is vital to square the probabilistic nature of QM with the proposition that quantum states apply to individual instances of systems.
+In terms of whether the textbooks only discuss the orthodoxy, or trivialise other interpretations, we find this has a >0.9>0.9 frequency pre-1980.
+This drops to $\sim 0.7$ for post-1980 but rises to $\sim 0.85$ for post-2000.
+Few rejections occur before 1980, with Weinberg [weinberg2013], Miller [miller2010], Omnes [Omnes1999-OMNUQM], Greiner [greiner1989], Rae [rae2002], and Fong [fong2013elementary] all providing more nuanced discussions of interpretation.
+Notably, the rejection rate drops significantly in the post-2000 sample (to below that of even the pre-1980 case).
+This odd result suggests more recent books talk about other interpretations less, even as the orthodoxy fragments (as we will see below).
+Perhaps this is typified by Griffiths [Griffiths2004Introduction], who views Bell inequality violations as empirically deciding in favour of the orthodoxy.
 Next, we can ask whether there are correlations between proposition responses.
@@
 We choose only to display those for which the assent frequency exceeds $0.6$ to determine if there is a strong correlation between orthodox positions.
-Importantly, the strongest correlation occurs for the two pillars of the orthodoxy: “measurement disturbs systems” and “non-ensemble $\psi$”, indicating that they are correlated 9595% of the time.
-Over all, it is evident that relatively strong pair-wise correlations exist between all the propositions with assent frequency above $0.6$.
-They all correlate strongly to one (or both) of the two pillar propositions as well.
+Importantly, the strongest correlation occurs for the pillars of the orthodoxy: “measurement disturbs systems”, “no pre-measurement values”, and “non-ensemble $\psi$”, being correlated $\gtrsim 80$% of the time.
+Over all, it is evident that relatively strong pair-wise correlations exist between almost all the propositions with assent frequency $\gtrsim 0.6$.
+They all correlate strongly to at least one of the pillar propositions as well.
 This demonstrates that selecting an orthodoxy on this basis yields a consistently affirmed set of propositions.
+The only exception here is “uncertainty linked to disturbance”, which generally features in all the weaker correlations.
+This isn’t too surprising, as it falls below a $0.6$ affirmation in post-1980 works.
 Proposition 1 Proposition 2 Full sample Post-1980
@@
 Collapse Single particle wave-like 0.79(0.76) 0.81(0.81)
-Collapse Measurement creates values 0.81(0.52) 0.81(0.56)
 Collapse Measurement disturbs system 0.81(0.76) 0.81(0.78)
+Collapse No pre-measurement values 0.81(0.76) 0.85(0.81)
+Collapse Orthodoxy only 0.71(0.62) 0.70(0.59)
 Non-ensemble $\psi$ Ontologically complete 0.83(0.81) 0.81(0.78)
@@
 Non-ensemble $\psi$ Single particle wave-like 0.86(0.83) 0.81(0.78)
-Non-ensemble $\psi$ Measurement creates values 0.90(0.57) 0.85(0.56)
 Non-ensemble $\psi$ Measurement disturbs system 0.95(0.90) 0.96(0.89)
+Non-ensemble $\psi$ No pre-measurement values 0.90(0.88) 0.89(0.85)
+Non-ensemble $\psi$ Orthodoxy only 0.83(0.74) 0.78(0.63)
 Ontologically complete Wave-particle duality 0.64(0.64) 0.63(0.63)
@@
 Ontologically complete Single particle wave-like 0.79(0.79) 0.78(0.78)
-Ontologically complete Measurement creates values 0.81(0.52) 0.78(0.48)
 Ontologically complete Measurement disturbs system 0.81(0.79) 0.81(0.78)
+Ontologically complete No pre-measurement values 0.79(0.76) 0.78(0.74)
+Ontologically complete Orthodoxy only 0.76(0.69) 0.74(0.63)
 Wave-particle duality Single particle uncertainty 0.67(0.67) 0.59(0.59)
 Wave-particle duality Single particle wave-like 0.76(0.76) 0.74(0.74)
-Wave-particle duality Measurement creates values 0.83(0.55) 0.85(0.63)
 Wave-particle duality Measurement disturbs system 0.74(0.71) 0.74(0.70)
+Wave-particle duality No pre-measurement values 0.76(0.74) 0.78(0.74)
+Wave-particle duality Orthodoxy only 0.64(0.57) 0.59(0.48)
 Single particle uncertainty Single particle wave-like 0.83(0.81) 0.78(0.74)
-Single particle uncertainty Measurement creates values 0.83(0.50) 0.74(0.44)
 Single particle uncertainty Measurement disturbs system 0.86(0.83) 0.81(0.78)
-Single particle wave-like Measurement creates values 0.86(0.55) 0.81(0.56)
+Single particle uncertainty No pre-measurement values 0.81(0.76) 0.74(0.67)
+Single particle uncertainty Orthodoxy only 0.79(0.74) 0.70(0.63)
 Single particle wave-like Measurement disturbs system 0.83(0.81) 0.81(0.78)
-Measurement creates values Measurement disturbs system 0.90(0.55) 0.89(0.56)
-Table 3: The frequency with which answers for two propositions match in the full and post-1980 samples. The values in brackets reflect the case without inference of responses.
+Single particle wave-like No pre-measurement values 0.86(0.83) 0.85(0.81)
+Single particle wave-like Orthodoxy only 0.79(0.74) 0.74(0.67)
+Measurement disturbs system No pre-measurement values 0.86(0.81) 0.85(0.78)
+Measurement disturbs system Orthodoxy only 0.81(0.74) 0.78(0.67)
+No pre-measurement values Orthodoxy only 0.79(0.69) 0.74(0.59)
+Table 4: The frequency with which answers for two propositions match in the full and post-1980 samples. Values in brackets indicate those without inferred responses.
 Finally, we have Figure 5. This displays the “orthodoxity” of each book, i.e. how close they are to the averaged responses over the full sample. This is determined following
@@
 where the sum runs over the propositions, $w_{i}$ is the average response value, and $a_{i}$ is that of the reference work in question.
-Note that we exclude propositions that do not have a full-sample assent frequency >0.6>0.6.
+Note that we exclude propositions that do not have a full-sample assent frequency >0.6>0.6 as well as “orthodoxy only”.
 As illustrated above, this selection criterion results in set of well-correlated propositions.
-The average orthodoxity is 7474% with standard deviation 2323% indicating a robust degree of agreement on average.
-This includes inferred values, but, this effect is small, going to 7171% and 2424% without them.
+The average orthodoxity is 6666% with standard deviation 2626% indicating there isn’t a very robust degree of orthodoxity on average.
+This includes inferred values, but, this effect is small.
 Note that these averages are robust to increasing the frequency inclusion threshold to $0.7$ as well.
@@
 This allows us to determine if there are any trends in the orthodoxy.
-Early works adhere very closely to the full-sample orthodoxy, with those before 1980 having an average orthodoxity of 8383% with standard deviation 1414%.
+Early works adhere very closely to the full-sample orthodoxy, with those before 1980 having an average orthodoxity of 7777% with standard deviation 1616%.
 The period around the early 90’s shows a fluctuation of far lower agreement with the orthodoxy.
-This results in post 1980 works having an average orthodoxity of 6868% and standard deviation 2525%.
-The substantial nature of this change from the earlier consensus is most vividly illustrated by the 50% increase in the post-1980 standard deviation, despite a 7373% increase in sample size.
-The 90’s dip in the orthodoxity is constituted by the works: Peres [38], Moore [36], Griffiths [37], Mandl [39], Rae [15], and Sakurai [42].
+This results in post 1980 works having an average orthodoxity of 6060% and standard deviation 2828%.
+The substantial nature of this change from the earlier consensus is most vividly illustrated by the post-1980 standard deviation almost doubling that of pre-1980, despite a 7373% increase in sample size.
+The 90’s dip in the orthodoxity is constituted by the works: Peres [peres1993], Moore [moore1998], Griffiths [Griffiths2004Introduction], Mandl [mandl2013quantum], Rae [rae2002], and Sakurai [Sakurai:1167961].
 Of these, Mandl is here mainly due to omission, rather than contesting orthodox positions.
-If we replace Mandl with Weinberg [21], we form a set of all the authors who deny more than one proposition.
+If we replace Mandl with Weinberg [weinberg2013], we form a set of all the authors who deny more than one proposition.
 However, there is no consistent pattern in these denials.
+Interestingly, the average orthodoxity rises to $66\pm 28$% as we transition to the post-2000 sample.
+Combined with the large drop-off in denials of “orthodoxy only”, this suggests there is some resurgent support for the orthodoxy.
+However, the standard deviation remaining nearly double the pre-1980 value indicates that fragmentation continues.
 Overall, there appears to be a post-1980 divergence from the earlier orthodoxy.
@@
 This suggests the possibility that our present epoch is the start of a collapse in confidence in the orthodoxy.
-A survey of working physicists has recently demonstrated that the community is fragmented on the interpretation of quantum mechanics [2], notably more so than a previous survey in 2016 [1].
+A survey of working physicists has recently demonstrated that the community is fragmented on the interpretation of quantum mechanics [survey2025], notably more so than a previous survey in 2016 [sivasundaram2016surveyingattitudesphysicistsconcerning].
 This could well signal an increase in divergence from the previous quantum orthodoxy in future textbooks.
 Somewhat complicating such an idea is the tendency of more orthodox post-1980 works to cite Bell’s theorem as a vindication of many aspects of the orthodoxy, perhaps accounting for the apparent resurgence in orthodoxity after the late 1990’s.
-The timeline notably aligns with the emergence of “Copenhagen+” interpretations like quantum Bayesianism (see [60]).
-For additional works that were not included in the sample we find the following orthodoxity measures: Ballentine [14] 0.88-0.88, Home [6] 0.42, Hannabuss [17] 0.45, and Nielsen & Chuang [18] 0.42.
-The last of these, Nielsen & Chuang [18], only shows a lower orthodoxity by virtue of never mentioning the topic of waves.
-Their strong endorsement of Peres [38] and Sakurai [42] might indicate little sympathy with this aspect of the orthodoxy, but the clarity of this is marred by also endorsing Cohen-Tannoudji [45].
-Interestingly, Home’s characterisation of the orthodoxy [6] appears incomplete at $0.42$.
+The timeline notably aligns with the emergence of “Copenhagen+” interpretations like quantum Bayesianism (see [timpson_quantum_2008]).
+For additional works that were not included in the sample we find the following orthodoxity measures: Ballentine [ballentine1998quantum] 0.88-0.88, Home [Home1997] 0.3, Hannabuss [Hannabuss:1999fa] 0.16, and Nielsen & Chuang [nielsen2010] 0.3.
+The last of these, Nielsen & Chuang [nielsen2010], only shows a lower orthodoxity by virtue of never mentioning the topic of waves.
+Their strong endorsement of Peres [peres1993] and Sakurai [Sakurai:1167961] might indicate little sympathy with this aspect of the orthodoxy, but the clarity of this is marred by also endorsing Cohen-Tannoudji [Cohen-Tannoudji:101367].
+Interestingly, Home’s characterisation of the orthodoxy [Home1997] appears incomplete at $0.3$.
 Additionally, this is only when one concentrates on the aspects they attributed to the non-ensemble case.
@@
 This principle as the central pillar of the orthodoxy is apriori dubious.
-This is because the notion of collapse is a supplement to the basic mathematical formalism [61, 14, 21], in the form of an additional physical/quasi-physical process with attendant projection postulate.
+This is because the notion of collapse is a supplement to the basic mathematical formalism [Wallace2019, ballentine1998quantum, weinberg2013], in the form of an additional physical/quasi-physical process with attendant projection postulate.
 Since Copenhagen-like interpretations hinge so heavily on collapse, their popularity implies that Home’s primary condition of orthodoxy becomes more doubtful.
-Interestingly, Ballentine [14], who would express an orthodox view according to Home, achieves an orthodoxity of 0.88-0.88 here, dissenting from (or not addressing) every proposition considered.
+Interestingly, Ballentine [ballentine1998quantum], who would express an orthodox view according to Home, achieves an orthodoxity of 0.88-0.88 here, dissenting from (or not addressing) every proposition considered.
 This is because Home’s rationale for including ensemble views in the orthodoxy may not be well justified.
@@
 Figure 5: The “orthodoxity” or degree to which each studied source agrees with the averaged responses to each proposition. This is ordered by date of first publication.
-We can study how time period effects the orthodoxy by change the orthodoxy weights to those of a subsample.
+We can study how time period effects the orthodoxy by changing the orthodoxity weights to those of a subsample.
 We display the changes in orthodoxy for both pre and post-1980 selections in Fig.
-4.
-Importantly, the pre-1980 case includes two extra propositions: “no trajectories” and “complementarity”, as both pass the $0.6$ assent threshold in that subsample.
-For both our modified orthodoxies, the mean changes from the full sample are small, being $0.0049$ and $0.015$ respectively.
-This indicates a small preference for the pre-1980 orthodoxy, with several authors becoming around 1010% more orthodox in this case.
-In contrast, the post-1980 case has smaller orthodoxity gains and several large decrements.
+6.
+Importantly, the pre-1980 case includes three extra propositions: “no trajectories”, “uncertainty linked to disturbance”, and “complementarity”, as all pass the $0.6$ assent threshold in that subsample only.
+For both our modified orthodoxies, the mean changes from the full sample are small, being $0.012$ and $0.006$ respectively.
+This indicates a small preference for the post-1980 orthodoxy, with several authors becoming around 1010% more orthodox in this case but very few reducing.
+In contrast, the pre-1980 case has many decrements, despite larger gains.
 Interestingly, the largest post-1980 decreases are all for post-1980 works.
 Whereas, the largest gains for a pre-1980 orthodoxy are from post-1980 works as well.
-This indicates that the most robust orthodoxy can be drawn from pre-1980 works’ positions, but there are definite hints of evolution.
+This indicates there are definite hints of evolution in the orthodoxy.
 However, there are no significant correlations between the changes (for any of our temporal subsamples), indicating that a new consistent orthodoxy is yet to emerge.
 Figure 6: The changes in “orthodoxity” that result if we use different samples to define the orthodoxy.
-We can now ask how far the revealed orthodoxy diverges from the idea that it can be defined entirely by the ontological primacy of the quantum state (as advocated by Home [6]).
+We can now ask how far the revealed orthodoxy diverges from the idea that it can be defined entirely by the ontological primacy of the quantum state (as advocated by Home [Home1997]).
 The first element that brings this into question is the “non-ensemble $\psi$” proposition.
@@
 The two main pillars that are empirically revealed are “measurement disturbs systems” and “states do not just describe ensembles”.
-Notably, the significance of consensus around ontological completeness has declined for modern works, as a result of some authors [32, 36] instead positing “statistical completeness” or openly suggesting a more general theory should be found [21].
+Notably, the significance of consensus around ontological completeness has declined for modern works, as a result of some authors [phillips2013introduction, moore1998] instead positing “statistical completeness” or openly suggesting a more general theory should be found [weinberg2013].
 6 Conclusion
-When we consider a sample of $42$ textbooks, aimed at both graduate and undergraduate students of quantum mechanics, we find a robust orthodoxy emerging.
-This is powerfully illustrated by the average orthodoxity of the sources being 7373% with a standard deviation of 2424%.
-Given that the studied sample spans nearly a century, this provides good evidence that many aspects of the consensus have remained unchanged.
-In post-1980 works, the average orthodoxity is reduced to 6868% from a pre-1980 value of 8484%.
-In particular, there is less support for the notions that individual particles are wave-like, and that the quantum state is ontologically complete.
+When we consider a sample of $42$ textbooks, aimed at both graduate and undergraduate students of quantum mechanics, we find evidence of robust orthodoxy that has begun to fragment.
+This is powerfully illustrated by post-1980 works having an average orthodoxity of 6060%, down from a pre-1980 value of 7777%.
+The standard deviation almost doubles from 1616% to 2828% when moving from pre to post-1980 samples.
+For the full sample, the average orthodoxity of the sources is 6666% with a standard deviation of 2525%.
+Over time, there has been less support for the notions that individual particles are wave-like, and that the quantum state is ontologically complete.
 There are also more significant reductions in support for the notions that the uncertainty principle applies to individual particles, quantum uncertainty results from measurement disturbance, particles do not have trajectories, and complementarity.
-The reduction in orthodoxy over time mirrors recent surveys of physicists attitudes in quantum mechanics [2].
+The reduction in orthodoxy over time mirrors recent surveys of physicists attitudes in quantum mechanics [survey2025].
 Perhaps indicating the start of opinion fragmenting away from the orthodoxy, as there are no obvious patterns for groups of propositions being denied together.
-The revealed orthodoxy strongly resembles many aspects of Copenhagen-like frameworks.
-Particularly, there is powerful agreement (assent frequency $0.91$) that measurement inherently disturbs quantum systems, single instances of systems are individually described by a quantum state, rather than as an ensemble (assent frequency $0.96$), state collapse happens (assent frequency $0.79$), that individual particles are wave-like (assent frequency $0.88$), and that quantum states are ontologically complete (assent frequency $0.85$).
-These positions are further supplemented by the existence of a wave-particle duality (assent frequency $0.78$) and that measurement actually creates or alters the definite property values being measured (assent frequency $0.76$, but only $0.59$ without inferred values).
-More contentious, but still popular, are the notions that the uncertainty principle arises from measurement disturbance (assent frequency $0.67$), and that individual particle measurements are subject to the uncertainty principle (assent frequency $0.85$).
-It is clear that the orthodoxy favours the view of the quantum state as a complete description of the actual state of affairs, rather than the more instrumental view attributed to Copenhagen by authors like Stapp [4].
-In this regard, the views of the orthodoxy espoused by Home [6] and Jammer [5] are vindicated.
-The work most embodying Stapp’s version would be Peres [38], with similarities in Mandl [39], Nielsen & Chuang [18], Hannabuss [17], and Sakurai [42].
+For the full sample of works, the revealed orthodoxy strongly resembles many aspects of Copenhagen-like frameworks.
+Particularly, there is powerful agreement (assent frequency $0.91$) that measurement inherently disturbs quantum systems, quantum systems have no pre-measurement values ($0.91$ assent), single instances of systems are individually described by a quantum state, rather than as an ensemble (assent frequency $0.96$), state collapse happens (assent frequency $0.79$), that individual particles are wave-like (assent frequency $0.88$), and that quantum states are ontologically complete (assent frequency $0.85$).
+These positions are further supplemented by the existence of a wave-particle duality (assent frequency $0.76$) and that individual particle measurements are subject to the uncertainty principle (assent frequency $0.85$).
+More contentious, but still popular, is the notion that the uncertainty principle arises from measurement disturbance (assent frequency $0.67$).
+It is clear that the orthodoxy favours the view of the quantum state as a complete description of the actual state of affairs, rather than the more instrumental view attributed to Copenhagen by authors like Stapp [stapp1972].
+In this regard, the views of the orthodoxy espoused by Home [Home1997] and Jammer [Jammer1974-JAMTPO-10] are vindicated.
+The work most embodying Stapp’s version would be Peres [peres1993], with similarities in Mandl [mandl2013quantum], Nielsen & Chuang [nielsen2010], Hannabuss [Hannabuss:1999fa], and Sakurai [Sakurai:1167961].
 However, it is not clear that all the logical consequences of Home’s view (that the orthodoxy is entirely characterised by ontological completeness) are believed by the physics community (and it is clear there are wave-related addenda).
@@
 In fact, it is a strong requirement of the orthodoxy that quantum states describe single instances, rather than ensembles.
-It is also interesting to note that respondents in [1] characterise the Copenhagen interpretation’s most salient features as “collapse” (77%), “complementarity” (71%), “indeterminism” (46%), and “correspondence” (43%).
+It is also interesting to note that respondents in [sivasundaram2016surveyingattitudesphysicistsconcerning] characterise the Copenhagen interpretation’s most salient features as “collapse” (77%), “complementarity” (71%), “indeterminism” (46%), and “correspondence” (43%).
 This is quite different to the orthodoxy deduced from our sample of textbooks.
-Where complementarity is far less important, indeterminism is more strongly supported (non-ensemble $\psi$ + ontological completeness), and collapse is popular but not absolutely primary.
+Where complementarity is far less important, indeterminism is more strongly supported (non-ensemble $\psi$ + no pre-measurement values), and collapse is popular but not absolutely primary.
 This suggests that, although physicists might predominantly report subscribing to the Copenhagen interpretation, the prevailing community beliefs do not fully conform to their own characterisation of this viewpoint.